### v1

# medR: Reward Engineering for Clinical Offline Reinforcement Learning via Tri-Drive Potential Functions

Abstract
Reinforcement Learning (RL) offers a powerful framework for optimizing dynamic treatment regimes (DTRs). However, clinical RL is fundamentally bottlenecked by reward engineering: the challenge of defining signals that safely and effectively guide policy learning in complex, sparse offline environments. Existing approaches often rely on manual heuristics that fail to generalize across diverse pathologies. To address this, we propose an automated pipeline leveraging Large Language Models (LLMs) for offline reward design and verification. We formulate the reward function using potential functions consisted of three core components: survival, confidence, and competence. We further introduce quantitative metrics to rigorously evaluate and select the optimal reward structure prior to deployment. By integrating LLM-driven domain knowledge, our framework automates the design of reward functions for specific diseases while significantly enhancing the performance of the resulting policies.
Machine Learning, ICML
1 Introduction
Critical care medicine is fundamentally defined by high-stakes, sequential decision-making, RL is a natural fit for optimizing DTRs where an agent must make sequential decisions that adapts to a patient’s changing state over time. Reward serves as the core of RL that not only guides the agent learning but also defines the optimization problem (Zhu et al., ). Small changes in reward function weights can cause profound policy changes(Roggeveen et al., 2024) and bad reward signals affect beyond “win-or-lose” in gaming and can lead to “life-or-death” in healthcare. Errors in sensitive fields such as autonomous driving have already been shown to be problematic (Knox et al., 2023). However, reward engineering in clinical RL is less studied and benchmarked. This underscores an urgent need to design better rewards that align with clinical goals (Basu et al., 2025).
Current rewards in RL for DTRs largely depend on manual and static design. Drawing an analogy to recent advances in LLM, we categorize existing clinical reward mechanisms into three distinct paradigms: sparse Outcome Reward Model (ORM), which assigns scalar rewards based on terminal patient outcomes like survival (Komorowski et al., 2018). The second is dense Process Reward Model (PRM), which depends on dense feedback based on intermediate physiological state changes (Zhang et al., 2024c). To mitigate the sparsity of terminal outcomes, the standard approach augments them with intermediate feedback to form a composite reward ($R_{OPRM}=R_{ORM}+R_{PRM}$) (Raghu et al., 2018; Lu et al., 2024). In this paradigm, clinicians manually select disease-relevant features and encode them using simple heuristic logic. However, the terminal reward is often weighted disproportionately high, overshadowing the dense shaping signals and the selection of specific parameters and weights remains largely ad-hoc, lacking systematic justification (Lu et al., 2024; Zhang et al., 2024c).
Despite efforts in designing various reward signals to accelerate policy learning, there are currently no guidelines on designing effective rewards for clinical RL due to several major challenges. First, outcomes in clinical settings are sparse and delayed. While mortality is the ultimate ground truth outcome, it occurs only at the end of ICU stay which leads to credit assignment problem. Second, the target in a treatment process is often implicit and complicated. It extends beyond binary survival to encompass the quality of recovery. There is a profound gap between the high-dimensional complexity of patient physiology and the scalar reward value required by RL algorithms. This gap leads to misaligned objectives, such as the “Pyrrhic Victory”(Livingston et al., 2019) where an agent optimizes for survival via aggressive interventions or reward hacking. Third, rewards are not verifiable. Unlike robotics or games, clinical RL is strictly offline, precluding online exploration to validate reward signals. Consequently, current rewards rely on brittle, manual heuristics that lack interpretability and fail to generalize across different diseases or datasets.
Recent advancements have identified LLMs as a promising solution to bridge the gap between high-level human intent and low-level reinforcement signals. While early work utilized LLMs as direct evaluators of agent behavior(Kwon et al., 2023), more recent methods focus on using LLMs to generate executable reward function codes(Ma et al., 2023; Xie et al., 2023). We leverage this generative capability to construct a reward function that captures the nuance of DTRs without the prohibitively high cost of manual trial-and-error engineering. Our pipeline leverages LLMs to bridge the semantic gap between medical knowledge and mathematical representation, utilizing a “Tri-Drive” mechanism comprising Survival, Confidence, and Competence to guide both reward generation and offline verification. Our specific contributions are:
  • We propose an automated pipeline for clinical RL that utilizes LLMs to perform interpretable feature selection and formulate a novel reward structure based on potential functions.
  • We design the reward functions via a Tri-Drive mechanism and derive corresponding offline quantitative metrics to act as a proxy for ground truth and verify reward functions.
  • We empirically validate our framework on three distinct, high-stakes clinical tasks and demonstrate that policies trained with our automated rewards consistently outperform those using standard rewards, achieving 77.3%, 66.7%, and 60.3% higher WIS scores over clinician baselines on three tasks, proving the framework’s generalizability across diverse diseases.
Figure 1: Framework pipeline of medR.
2 Background
We model the underlying clinical dynamics as a Partially Observable Markov Decision Process (POMDP), defined by the tuple $(\mathcal{S}^{*},\mathcal{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma)$. Here, $\mathcal{S}^{*}$ denotes the unobservable, true physiological state of the patient, and $\mathcal{A}$ represents the space of clinical interventions. The agent can not access $\mathcal{S}^{*}$ directly; instead, it receives intermittent observations $o_{t}\in\Omega$ like vital signs and lab results $\mathcal{O}(o|s^{*},a)$. To ensure computational tractability for offline policy learning, we approximate this POMDP as a Markov Decision Process (MDP) defined by $(\mathcal{S},\mathcal{A},P,R,\gamma)$ during policy training. We explicitly construct the proxy state space $\mathcal{S}$ in our reward functions to capture temporal irregularities inherent in clinical data: a state is defined as $s_{t}=(o_{t},\Delta t)$, where $\Delta t$ represents the time elapsed since the last observation. While standard MDP approximations often discard the uncertainty arising from partial observability, we explicitly mitigate this gap through our reward engineering framework.
2.1 Intrinsic Motivation
In this work, our reward function design follows the evolutionary perspective proposed in (Singh et al., 2009). They distinguish between reward signals and fitness functions where reward signals are internal scalar values that drive learning within an agent and fitness is the external evaluation that measures an agent’s success. For bounded agents in an offline environment, the optimal reward function is not the fitness function(Singh et al., 2010). The design of reward functions should optimize intrinsic motivations (IM), unlike previous literature about IM that focuses on incentivizing exploration which is dangerous in our offline setting, in our paper, we form it as a threefold drive: the drive to restore homeostasis(survival), the drive to lower uncertainty(confidence), and the drive to minimize cost(competence). In Hull’s Drive Reduction Theory(Hull, 1945), organisms are motivated to maintain physiological equilibrium, we treat homeostasis as a proxy of survival to optimize in the reward function. Unlike curiosity-driven frameworks that encourage exploration under uncertainty(Berlyne, 1960; Schmidhuber, 1991), in order to design a reward function specific to our offline environment(Singh et al., 2009), we conversely suppress uncertainty due to the sensitivity in critical care(Djulbegovic, 2011). White (1959) furthers the learning process of effective interaction with the environment, which is transferrable as effective treatment under clinical setting.
2.2 Potential-Based Reward
[…]
3 Methodology
3.1 Interpretable Feature Selection
We leverage the semantic reasoning capabilities of LLMs to select critical features for constructing the reward functions. Figure1 shows the framework of our method. We first compute a statistical summary for all candidate features in the dataset $\mathcal{D}$. Let $f$ be a candidate feature and $y$ be the patient outcome mortality. We generate a metadata tuple $\mathcal{M}_{f}=(\mu_{f},\sigma_{f},\eta_{f},\rho_{f}^{outcome},\rho_{f}^{action})$, where $\mu$ is the mean and $\sigma$ is the standard deviation of the non-missing values. We explicitly analyze the missingness rate $\eta$ because Electronic Health Records(EHRs) are inherently sparse, simply using imputed data for reward computation is misleading and can introduce bias. $\rho_{outcome},\rho_{action}$ are the correlation between feature $x$ and patient outcome $y$ or action $a$, for each mortality and action dimension, we compute the Pearson correlation coefficient with every feature $f$ in the state space.
We create a structured prompt that feeds this statistical metadata $\mathcal{M}$ into an LLM. The LLM acts as a clinical data scientist to select $K$ critical state features that affect the specific disease treatment. Crucially, for every selected feature, the model must generate a rationale explaining its selection. To mitigate the hallucination and instability inherent in LLM generation, we employ a stochastic ensemble selection strategy. The final feature set $\mathcal{F}^{*}$ is determined via majority voting: $\mathcal{F}^{*}=\{f\in\mathcal{F}_{all}\mid\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(f\in\mathcal{S}_{i})\geq\tau\}$ where $\mathcal{S}_{i}$ is the set of features selected in iteration $i$, and $\kappa$ is the consensus threshold. We then double check with clinicians on the features chosen by LLM and the rationales behind them. In this way we effectively simulate the knowledge of multiple clinicians to vote on the most robust candidates and validate with real clinicians to guarantee feature reliability.
3.2 Potential Based Reward Function Generation
We design a health potential that reflects the quality of physiological states. Past reward functions often define specific rules that directly scores the action, for example, Reward +1+1 if $70<mbp<80$, which will likely cause reward hacking and encouraging actions that can directly increase blood pressure. Instead, we form it as the gradient of potential to capture the patient state change. To bridge the gap between sparse outcome labels and dense clinical supervision, the reward function $R$ is composed of three distinct drives—Survival, Confidence, and Competence leveraging the intrinsic motivation theory (Singh et al., 2009). We define the reward function as follows:
$\small R(s,a,t,s^{\prime},a^{\prime},t^{\prime})=\left[\gamma\Phi(s_{t+1},t+1)-\Phi(s_{t},t)\right]-\lambda\mathcal{C}(a_{t})$ (1)
$\small\Phi(s_{t},t)=\delta(t)\cdot\sum_{f\in\mathcal{F}}\omega_{f}\cdot\mathcal{S}_{f}(v_{t,f})\cdot\mathcal{U}_{f}(\Delta t_{t,f})$ (2)
where $\gamma$ is the discount factor, $\Phi(\cdot)$ is the health potential, $\delta(t)$ is the strategic decay, $\mathcal{S}_{k}(v_{t,f})$ is the survival score with confidence weight $\mathcal{U}_{f}(\Delta t_{t,f})$ and $\mathcal{C}(\cdot)$ is the competence cost.
We use physiological stabilization as the proxy for survival. By embedding the metadata of critical features as part of the prompt, we personalize the reward design to handle critical patients with comorbidities. The LLM is prompted to design a potential function that rewards states closer to homeostasis. In the offline clinical setting, data irregularity poses a significant risk: a normal heart rate recorded 12 hours ago is not a reliable indicator of current stability. Since we are unable to refresh the data by ordering new lab tests, we utilize the elapsed time $\Delta t$ as indicators of freshness of the data $s_{t}=(o_{t},\Delta t_{t})$. The LLM generates confidence weights for each critical feature that inversely scales the reward based on $\Delta_{t,f}$. This mechanism penalizes stale states, effectively discounting the survival reward when data is uncertain to reduce epistemic uncertainty. We also add a strategic decay to prevent unnecessarily longer ICU stay. To prevent the agent from achieving survival via excessive or harmful interventions, we incorporate a competence component focused on treatment efficiency. Mathematically, this serves as a regularization term: if two treatments result in the same physiological transition $s_{t}\rightarrow s_{t+1}$, the potential function assigns a higher value to the less invasive action, penalizing over-treatment or flailing behavior. By decoupling the accumulating action cost $\mathcal{C}(a)$ from the telescoping potential $\Phi(s,t)$D.1, we effectively formulate the RL objective as the Lagrangian of a Constrained Markov Decision Process (CMDP). The agent maximizes the Lagrangian $\mathcal{L}=\text{Physiology}-\lambda\cdot\text{Toxicity}$, ensuring that high-stability states are only pursued if the cost of the intervention path does not exceed the value of the stabilizationD.2. Instead of manual heuristic tuning, the LLM acts as the logic designer. We feed the definitions of these three drives along with the feature metadata into the LLM, which autonomously synthesizes the logic into executable Python code for the reward function. To mitigate the risk of hallucination and capture diverse clinical perspectives, we generate $N$ distinct candidate functions in parallel, creating a diverse pool of candidates for the subsequent verification phase.
3.3 Tri-Drive Reward Selection
Since online verification is ethically prohibitive in clinical settings, we cannot evaluate reward candidates by deploying them on patients. Instead, we propose a set of offline fitness metrics to rigorously assess the quality of each generated reward function $R$. We formulate reward design as a Tri-Drive Selection Problem aiming to balance three conflicting objectives based on trajectory-level labeled ground truth: Survival, Confidence, and Competence Fitness.Survival Fitness ($J_{surv}$) validates whether the induced reward signal correlates with actual patient outcomes. We define a ground-truth trajectory score $G(\tau)$ that consists of both mortality and Sequential Organ Failure Assessment(SOFA) score(Lambden et al., 2019) increase compared to baseline(time when admitted to ICU). The fitness is the correlation $\rho$ between the cumulative reward and $G(\tau)$:
$G(\tau)=\mathbf{1}(\text{Survival})+\frac{1}{T}\sum_{t=0}^{T-1}\mathbf{1}(|\text{SOFA}_{t}-\text{SOFA}_{baseline}|<\epsilon)$ (3)
$J_{surv}(R)=\rho\left(R_{\tau},G(\tau)\right)$ (4)
Confidence Fitness ($J_{conf}$) penalizes reward functions that assign high credit to decisions made under high epistemic uncertainty. We define an uncertainty proxy $U(\tau)$ as the average time-gap $\Delta t$ of all the features $f$ in the critical feature set $F$. The fitness is the negative correlation between the cumulative reward and trajectory uncertainty:
$U(\tau)=\frac{1}{T\cdot F}\sum_{t=0}^{T-1}\sum_{f=1}^{F}\Delta t_{t,f}.$ (5)
$J_{conf}(R)=-\rho\left(R_{\tau},U(\tau)\right)$ (6)
Competence Fitness ($J_{comp}$) ensures that the reward function prioritizes efficient strategies that achieve or maintain high physiological stability with minimal necessary intervention. We define feature type as $\mathcal{T}=\{N,D_{l},D_{h}\}$ where $N$ denotes Normal-Range features, and $D_{l},D_{h}$ denote Directional features where minimization or maximization is physiologically optimal, respectively. We define the unified homeostasis score $H(x)\in[0,1]$ for a feature of type $T\in\mathcal{T}$ as:
$\small h(f)=\begin{cases}\mathbf{1}_{f\in\mathcal{I}}+\mathbf{1}_{f\notin\mathcal{I}}\cdot\sigma\left(k\left[0.5-\frac{d(f,\mathcal{I})}{\text{IQR}}\right]\right)&\text{if }T=N,\\ \sigma\left(k\left[0.5-f\right]\right)&\text{if }T=D_{l},\\ \sigma\left(-k\left[0.5-f\right]\right)&\text{if }T=D_{h},\end{cases}$ (7)
$H(s_{t})=\frac{1}{|F|}\sum_{f\in F}h(_{t,f}).$ (8)
We then define the Efficiency $E_{t}$ as the homeostasis increase penalized by the magnitude of the dose taken:
$E_{t}=H(s_{t+1})-H(s_{t})-\alpha\cdot\bar{a}_{t}.$ (9)
$J_{comp}=\rho\left(R_{\tau},E_{\tau}\right),$ (10)
where $\sigma(z)=(1+e^{-z})^{-1}$ is the standard sigmoid function, $d(f,\mathcal{I})$ is the distance from feature $f$ to the interval $\mathcal{I}$, IQR is the interquartile range width used for normalization. $\alpha$ and $k$ are cost parameters, $\bar{a}_{t}$ is the magnitude of the intervention. We employ the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) (Deb et al., 2002) to rank the generated candidates based on these three objectives. Instead of selecting a single weighted average, we identify the Pareto Frontier where no objective can be improved without degrading another. From this frontier, the solution closest to the utopia point representing the best trade-off.
4 Experiments
4.1 Tasks
We evaluate our framework on three distinct critical care tasks, covering different datasets, demographics, and action space types.
Sepsis Treatment We utilize the Medical Information Mart for Intensive Care (MIMIC)-IV v3.1(Johnson et al., 2020) database to learn optimal Intravenous (IV) fluid and Vasopressor administration strategies for sepsis patients. We select adult patients (age $\geq$ 18) satisfying the Sepsis-3 criteria (Singer et al., 2016), and exclude patients with missing mortality, demographics, and interventions, extreme missingness (>>60% missing data) or short hospital stays (<<24 hours). We define a septic shock starting from 24 hours before diagnosis and ending at 48 hours post-diagnosis. After filtration, 7,501 patients with 46 features are retained in the dataset. The action space is discrete consisting of two primary interventions: IV fluids and Vasopressors. We discretize IV fluid volume and vasopressor dosage into 5 bins each, resulting in a combined action space of $5\times 5=25$ discrete actions.
Mechanical Ventilation We use the multi-center eICU Collaborative Research Database (eICU-CRD)(Pollard et al., 2018) to optimize ventilator settings for patients with respiratory failure. The cohort includes adult patients who underwent invasive mechanical ventilation (MV) for at least 24 hours. We select patients aged $\geq$16 who received MV for at least 24 hours. We also exclude patients with missing mortality outcomes, demographics, interventions, and those ventilated for >>14 days. The final cohort contains 2168 patients with 41 features. The action space is discrete, controlling three key ventilator parameters: Positive End-Expiratory Pressure (PEEP), Fraction of Inspired Oxygen (FiO2), and Tidal Volume. We define a discrete action set based on clinically relevant adjustments for each parameter, resulting in a composite discrete action space of 18 actions (PEEP: 2 levels, FiO2: 3 levels, Tidal Volume: 3 levels).
Renal Replacement Therapy We apply our method to the AmsterdamUMCdb (Thoral et al., 2021) to optimize continuous renal replacement therapy (RRT) dosing. We select adult patients (age $\geq$ 18) who received CRRT during the ICU stay where the stay is no longer than 60 days. There are 857 patients with 18 features in the cohort. Unlike the previous tasks, the action space here is continuous. The agent controls the effluent flow rate clipped to the range $[0,60]$ ml/kg/h, a critical parameter determining the intensity of dialysis.
For all tasks, clinical time-series data (vitals, labs, medications) are aggregated into 1-hour time steps. Missing values are imputed using forward-filling for time-varying features. For CRRT effluent rate, we imputed values based on RRT sessions, defining a new session whenever the gap between two dialysis records exceeded 8 hours. All continuous state features are normalized to zero mean and unit variance to ensure training stability. For policy training, we use common safe offline RL methods, Batch-Constrained Q-Learning (BCQ)(Fujimoto et al., 2019) for discrete action space tasks (Sepsis and Ventilation), and Implicit Q-Learning (IQL)(Kostrikov et al., 2021) for continuous action space task (RRT).
4.2 Baselines
To rigorously evaluate the effectiveness of our automated reward design framework, we compare it against two categories of baselines: manually engineered heuristic rewards and direct LLM approaches.
4.2.1 Heuristic Reward Design
We implement three standard reward structures commonly used in clinical RL literature. These serve as the ”human-expert” baselines, representing varying degrees of clinical foresight:
Outcome Reward Model (ORM): A purely outcome-oriented function that assigns a non-zero reward only at the terminal step, +100+100 for survival (discharge) and 100-100 for mortality. The value of $\pm 100$ is used in both Sepsis(Komorowski et al., 2018) and Ventilation(Peine et al., 2021) tasks in previous literature. Since there is no past work that utilized outcome reward for RRT task, we adopt the same setting of $\pm 100$ for RRT. This baseline tests the agent’s ability to solve the credit assignment problem without intermediate guidance. We explicitly utilized ICU-discharge mortality rather than 90-day mortality as our terminal state definition in the baselines. We argue that an RL agent can only be held accountable for outcomes within its scope of control. An agent penalized for a death occurring 60 days post-discharge may erroneously learn that stabilizing and discharging a patient is suboptimal, leading to conservative strategies.
Process Reward Model (PRM): A process-oriented function that provides immediate feedback based on step-wise improvements in physiological stability. It incentivizes short-term stabilization without long-term survival awareness. We use the reward function fully based on SOFA score(Zhang et al., 2024c) for Sepsis. For Ventilation, though we note that there is one process reward based on ventilator free days(Yousuf et al., 2025), the reward calculated on our cohort is not dense enough for efficient policy training, we therefore adopt the intermediate reward from the OPRM(Liu et al., 2024). This further indicates that a fixed reward function may not be transferrable between different cohorts even with the same dataset. For RRT, we use the reward function based on Estimated Glomerular Filtration Rate (eGFR)(Zhang et al., 2024a).
Outcome Process Reward Model(OPRM): A weighted combination of the sparse and dense signals. This represents the state-of-the-art manual design, aiming to balance immediate physiological regularization with the ultimate goal of patient survival. We select from current SOTA methods (Perera et al., 2026)for Sepsis and (Liu et al., 2024)for Ventilation. Due to the lack of work on using OPRM for RRT, we choose to combine the PRM(eGFR) with ORM($\pm 100$) as OPRM.
4.2.2 LLM-based Design
We also compare against methods that leverage LLMs for reward generation without our proposed iterative refinement process. Since there are limited methods that leverage LLM to generate rewards for offline RL, we design two variants:
LLM-as-Reward (LLMR): In this setting, the LLM is used as a direct proxy for the reward function or a scorer of the clinicians’ actions. We embed the whole patient trajectory $(s,a,s^{\prime})$ with final outcome mortality as part of the prompt along with clinical context, based on the mortality of the patient, we let the LLM do credit assignment of +15+15 for survival or 15-15 for mortality along the trajectory. The LLM will choose steps that it thinks is the most crucial for the decision-making and the rest of the steps will be assigned 0. After the reward generation we will make sure the sum of reward is +15+15 or 15-15. This tests the LLM’s innate ability to evaluate clinical actions without explicit function coding.
Code Generation (CodeGen): The LLM is prompted to write a Python reward function in a single pass, given only the task description and list of available features without any further instructions on the reward components. This baseline evaluates the code-generation capability of the LLM absent the Tri-Drive and potential functions utilized in our method.
4.3 Evaluation Metrics
We employ a comprehensive set of metrics to evaluate both the intrinsic quality of the generated reward functions and the extrinsic performance of the resulting RL policies.
Action Agreement among Survivors: We calculate the percentage of time steps among the whole trajectory where the RL agent’s chosen action matches the clinician’s action, specifically within the cohort of patients who survived. High agreement on survivors suggests the agent learns safe, effective behaviors from successful clinical examples.
Off-Policy Evaluation(OPE) To quantify the policy’s theoretical performance, we employ Weighted Importance Sampling (WIS), a standard unbiased estimator for off-policy value estimation. To facilitate direct comparison, we use the same rewards generated by DeepSeek-R1 that was not involved in training nor as a baseline for WIS calculation on all the policies. We also plot WIS training curve to make sure that the return is higher as we train the policy.
Mortality vs. Cumulative Reward: To prove that our reward is highly correlated with ground truth mortality, we plot the observed patient mortality rate against cumulative reward along patient trajectories assigned by the clinician’s policy. A monotonic decreasing trend indicates that the learned value function correctly identifies patient risk. We compare with PRM which is the only baseline that also does not contain mortality as part of rewards.
Table 1: Agreement rates among survivors compared to clinician actions across tasks.
Method Sepsis Ventilation RRT
Joint IV Fluids Vasopressors Joint PEEP FiO2 Tidal Volume Dose On/Off
ORM $25.37\pm 0.78$ $30.01\pm 0.69$ $77.98\pm 1.01$ $19.79\pm 3.13$ $74.80\pm 3.13$ $47.87\pm 2.26$ $48.23\pm 3.41$ $50.77\pm 15.96$ $67.54\pm 16.14$
PRM $27.54\pm 1.03$ $33.17\pm 0.96$ $78.47\pm 0.95$ $21.24\pm 2.37$ $74.49\pm 3.26$ $49.34\pm 2.60$ $50.70\pm 3.51$ $54.64\pm 10.51$ $74.45\pm 3.03$
OPRM $27.51\pm 1.39$ $33.04\pm 1.27$ $78.44\pm 1.08$ $21.25\pm 2.58$ $73.24\pm 3.63$ $49.30\pm 2.22$ $50.02\pm 2.88$ $56.46\pm 11.66$ $74.83\pm 2.46$
[…]
4.4 RQ1: Should we use mortality as fitness or reward?
A prevalent assumption in clinical RL is that agents trained solely on sparse mortality outcomes will asymptotically converge to optimal behavior via automatic temporal credit assignment. However, this premise rarely holds true in the offline, high-noise, stochastic regime of critical care, and can lead to “superstitious learning.” In ICU a robust patient may survive despite suboptimal actions, while a high-severity patient may die despite optimal care. As shown in Table1 and Table8, the policy trained under ORM performs significantly worse than the other baseline and our method on the agreement among survivors which reveals that the former struggles to distinguish between policy quality and patient stochasticity. Simply maximizing survival may move the patient away from wellbeing. In Table 2, the ORM reward also results in the lowest return in both Sepsis and Ventilation tasks. Furthermore, utilizing outcome that happens later can introduce hindsight bias, the training signal should only depend on information available at decision time. We avoid this by aligning the reward with the clinician’s reality, optimizing immediate physiology as a proxy of survival. We separate mortality from reward and instead use the hindsight ground truth as fitness to evaluate whether the reward function induces high-quality policies. This confirms that explicitly rewarding the process of stabilization provides a more reliable learning signal than the noisy outcome.
Figure 2: The relationship between cumulative process reward and patient mortality probability for different tasks. A valid process should demonstrate clear downward trend with mortality, baselines that include mortality as part of the reward design are excluded for fair comparison.
4.5 RQ2: Can we use naive shaping of outcome reward model?
To facilitate policy training previous literature actively utilizes clinically-guided dense rewards that include both terminal and intermediate rewards to form an OPRM(Raghu et al., 2017; Lu et al., 2024), which can be approximated as a naive reward shaping in a broader sense. As shown in Figure2, the standalone process reward is not well correlated with mortality, which is dangerous as a secondary outcome to guide policy learning. In Figure4, naive shaping candidates often dominate on the $J_{surv}$ axis but collapse on the $J_{comp}$ axis. This indicates a reward hacking phenomenon where the agent learns to accumulate stability bonuses by maintaining the patient in a safe but non-improving state, rather than efficiently steering them toward discharge. This is proved by Table2 where it achieves low WIS in both Ventilation and RRT tasks. In contrast, medR maintains high performance across all three objectives. Our method avoids this failure mode through the time-dependent decay factor $\delta(t)$ explicitly penalizes loitering. As $\Phi(s,t)$ decays towards zero over time, the agent is mathematically forced to achieve stabilization targets early in the episode to maximize the return. By utilizing the potential difference $\gamma\Phi_{t+1}-\Phi_{t}$, our method mathematically ensures that rewards are only generated by state improvement rather than state maintenance, effectively preventing the infinite horizon hacking observed in the baseline.
Table 2: Comparison of WIS scores with bootstrapped confidence intervals (95%CI).
Method Sepsis Ventilation RRT
Clinician (Baseline) 54.50-54.50 (55.49,53.51)(-55.49,-53.51) 56.94-56.94 (57.29,56.59)(-57.29,-56.59) 87.61-87.61 (92.51,82.83)(-92.51,-82.83)
ORM 45.34-45.34 (46.53,44.02)(-46.53,-44.02) 31.06-31.06 (32.90,29.20)(-32.90,-29.20) 34.95-34.95 (51.31,17.35)(-51.31,-17.35)
PRM 13.12-13.12 (14.03,12.24)(-14.03,-12.24) 24.97-24.97 (26.47,23.59)(-26.47,-23.59) 38.97-38.97 (51.82,26.13)(-51.82,-26.13)
OPRM 17.73-17.73 (18.93,16.63)(-18.93,-16.63) 24.30-24.30 (25.59,23.20)(-25.59,-23.20) 35.57-35.57 (51.75,19.90)(-51.75,-19.90)
[…]
4.6 RQ3: Do we need to include action as part of reward?
Previous reward function design rarely includes action as part of reward signals. Without action costs, clinical agents are prone to masking physiological deterioration with high-intensity interventions. While this improves the immediate state representation, it often induces long-term toxicity. In Figure5,6,7, policies trained by heuristic reward tend to give higher dosages compared to medR. By strictly penalizing the intervention magnitude, our full Tri-Drive objective forces the agent to solve a Lagrangian relaxation of the clinical problem: maximize stability subject to minimum necessary harm. Our policy achieves the highest $J_{comp}$ in all three tasks (Figure4) while achieving the highest WIS(Figure2).
4.7 RQ4: Do LLMs qualify as reward functions?
LLMs are rapidly developing and demonstrating excellent reasoning capabilities. Previous research has explored the possibility of directly using LLMs as planners for DTRs(Luo and Zhu, 2025) and concludes that they exhibit notable limitations. Therefore, a natural question is whether they can instead directly serve as the reward signal instead and output a scalar quality score. Our experiments indicate that LLMs are ill-suited for this direct role in the clinical offline RL setting. While an LLM can accurately identify immediate physiological abnormalities, it fails to perform temporal credit assignment. The standard LLMR baseline consistently demonstrates lower agreement with clinical policies compared to medR across all tasks, particularly in complex interventions like RRT dosing where it trails by over 15 percentage points ($50.63\%$ vs. $66.44\%$). Furthermore, LLMR exhibits notably higher variance, suggesting that generic LLM-generated rewards lack the stability and precision of our potential-based approach. While they naturally demonstrate comparably higher $J_{surv}$ since we force the sum of reward to be a direct indicator of mortality, they systematically collapse on $J_{conf}$ and $J_{comp}$, suggesting it generates survival-at-all-costs policies that ignore clinical efficiency and data uncertainty. By prompting the LLM with the full trajectory and outcome, the hindsight bias appears again and the model will tend to hallucinate positive or negative rewards along the whole trajectory. This confirms that while LLMs possess domain knowledge, they function as semantic pattern matchers rather than value function approximators.
Table 3: Ablation study. We evaluate the contribution of feature selection, the potential term, and Tri-Drive logic and prove medR demonstrates the most consistent performance across tasks.
Ablation Variant $J_{\text{surv}}$ $J_{\text{conf}}$ $J_{\text{comp}}$
Sepsis
w/o Feature Selection 0.61 0.47 0.55
w/o Potential 0.44 0.50 0.65
w/o Tri-Drive Logic 0.56 0.54 0.56
medR 0.68 0.57 0.69
Ventilation
w/o Feature Selection 0.65 0.77 0.55
w/o Potential 0.41 0.24 0.49
w/o Tri-Drive Logic 0.60 0.74 0.59
medR 0.69 0.82 0.61
RRT
w/o Feature Selection 0.50 0.46 0.42
w/o Potential 0.54 0.32 0.48
w/o Tri-Drive Logic 0.52 0.70 0.58
medR 0.60 0.66 0.60
4.8 RQ5: Can LLMs directly generate reward code?
[…]
Figure 3: WIS training plot.
4.9 Ablation Study
To rigorously evaluate the individual contributions of our framework’s components, we conducted an ablation study by systematically removing (1) the LLM-driven Feature Selection, (2) the Potential-based reward formulation, and (3) the Tri-Drive structured logic. Table 3 summarizes the impact of these components on policy fitness ($J_{\text{surv}}$, $J_{\text{conf}}$, $J_{\text{comp}}$) across all three clinical domains. The most critical component of our framework is the potential difference formulation ($\gamma\Phi_{t+1}-\Phi_{t}$). Removing this term results in a catastrophic collapse in performance, particularly in the Mechanical Ventilation domain, where $J_{\text{conf}}$ drops precipitously from $0.82$ to $0.24$. This confirms that without the telescoping sum property, the reward signal fails to distinguish between state maintenance and state improvement, leading to policies that are highly uncertain and poorly correlated with patient survival. The ablation of automated feature selection highlights the necessity of filtering the high-dimensional clinical state space. This effect is most pronounced in the RRT task, where removing feature selection degrades survival fitness ($J_{\text{surv}}$) from $0.60$ to $0.40$. In complex domains like renal replacement, irrelevant features introduce significant noise; our LLM-driven selection effectively isolates the causal variables required for stable learning. Finally, we evaluate the structural scaffolding of the Survival, Confidence, and Competence drives. While this variant occasionally achieves high scores in isolated metrics, for example, achieving the highest $J_{\text{conf}}$ of $0.70$ in RRT, it fails to maintain the Pareto balance required for effective care. For instance, in RRT, the increase in confidence comes at the cost of survival fitness. The full medR framework forces a necessary trade-off, optimizing for policies that are simultaneously safe, effective, and efficient.
5 Conclusion
In this work, we introduced medR, a novel framework for automating reward engineering in offline clinical RL. By leveraging LLMs to define Tri-Drive potential functions, medR successfully addressed policy optimization across three distinct clinical tasks. The substantial performance gains observed in both traditional OPE metrics and our proposed Tri-Drive fitness scores validate the efficacy of utilizing LLMs for offline reward shaping. Given the clinical necessity for reward functions that are personalized to specific diseases and patient cohorts, medR establishes a scalable foundation for applying reinforcement learning to broader medical challenges and more complex intervention spaces.
6 Impact Statement
This work addresses one of the most significant barriers to the deployment of RL in healthcare: the difficulty of defining mathematically precise reward functions that align with complex clinical objectives. By automating this process through medR, we lower the technical barrier for clinicians and researchers to develop custom treatment policies for diverse patient cohorts and rare pathologies, potentially democratizing access to precision medicine tools. Methodologically, our Tri-Drive framework contributes to AI safety by explicitly embedding uncertainty quantification (Confidence) and dosage constraints (Competence) into the reward structure, mitigating the risks of aggressive policy behavior common in offline RL. However, we acknowledge that LLM-generated rewards may inherit biases present in training data or prompt phrasing. While our results demonstrate high correlation with clinical outcomes, rigorous prospective validation and human-in-the-loop oversight remain essential prerequisites before any clinical deployment.
Appendix A Prompts
Prompt 1: Prompt for Feature Selection
You are an expert Clinical Data Scientist and Intensivist specializing in Offline Reinforcement Learning.
TASK: Select the top 7 **Critical State Features** based on the statistical analysis provided below.
SELECTION CRITERIA:
- Select features that are strong indicators of patient condition that are highly correlated to the disease and patient outcome.
- Exclude features that are direct proxies of interventions to prevent reward hacking (high correlation with actions).
- EXCLUDE all demographic and baseline features: age, gender, elixhauser vanwalraven, weight, readmission, step id. These are static patient characteristics, not dynamic state features. Focus ONLY on dynamic physiological and clinical state features that change over time.
- Prefer features with low missingness and strong predictive power for outcomes.
OUTPUT FORMAT: IMPORTANT: Output ONLY valid JSON. Do not include any preamble, thinking process, explanations, or text before or after the JSON.
Return a JSON object with keys: critical state features, rank the features by their importance for reward modeling. For each feature, provide a 1-sentence rationale explaining its selection based on the provided statistics.
Format your response as structured JSON with the following schema:
{"critical_state_features": [{
"feature_name": "...",
"rationale": "..."
Your response must start with " and end with ". Do not include any other text.
DATASET SUMMARY:
Total records:
Total patients:
Average records per patient:
Mortality Rates:
FEATURE STATISTICS:
Feature:
Count:
Mean:
Range:
Median: IQR:
CORRELATIONS WITH OUTCOMES:
Outcome: mortality
- feature\_name: r= (p=, n=)
ACTION-FEATURE CORRELATIONS (to identify action-dependent features):
Action: action1
- feature\_name: r= (p=, n=)
Prompt 2: Prompt for Reward Generation
[…]
You must **invent** the mathematical logic for three components based on the data below.
## PART 1: MATHEMATICAL DESIGN (You must think first)
The potential function consists of three components, survival, confidence, and competence. For each component, explicitly state the formula you will use.
1. **Survival:** Choose a function that scores physiology (0-1).
- *Design Constraint:* Use both clinical knowledge and feature statistics to define the healthy range. Higher potential when the physiological states move towards healthier range. "Goldilocks" features need Bell curves. Directional features need decay curves.
[…]
- *Design Constraint:* Trust must drop as time gap increases.
3. **Competence:** Choose a penalty function for actions.
- *Design Constraint:* Higher dose = Lower potential.
Also design a time decay funtion for strategy annealing, which decays the potential toward zero as the episode progresses. This is different from the confidence component.
### PART 2: PYTHON IMPLEMENTATION RULES (CRITICAL)
1. **NO DUMMY FUNCTIONS.** Every function you call must be defined in this script. Do not write _compute_survival()‘ unless you write the code for it immediately above.
2. **NO EXTERNAL CONFIG.** All dictionaries (Targets, Sigmas, Taus) must be defined as variables inside the script.
3. **BALANCED COPONENTS.** All component scores must be balanced and normalized, no single component overpowers others.
### INPUT FORMAT:
- state={’feature1’: (value1, delta1),
feature2’: (value2, delta2),
- Values for the features are normalized with min max values to the range of 0 to 1
[…]
- action={’action1’: value1,
action2’: value2,
- Values for the actions are in levels represented with integers (e.g. 0, 1, 2, 3 from low to high)
### OUTPUT FORMAT
Provide a single Python code block that defines the potential function and the reward function.
**CODE TEMPLATE (You MUST follow this structure):**
‘‘‘python
import numpy as np
import math
# --- 1. PARAMETER DEFINITIONS (HARD-CODED FROM DATA) ---
# Design your targets here based on the JSON summary
SURVIVAL_CONFIG = {
# ... FILL ALL FEATURES ...
CONFIDENCE_TAU = {
# ... FILL ALL FEATURES ...
# --- 2. MATH HELPER FUNCTIONS (NO PLACEHOLDERS) ---
def time_decay(t):
# WRITE THE TIME DECAY FUNCTION FOR STRATEGIC ANNEALING
return ...
def compute_survival_score(val, params):
# WRITE THE ACTUAL MATH HERE
return ...
def compute_confidence_weight(delta_t, tau):
# WRITE THE ACTUAL EXPONENTIAL DECAY MATH
return ...
def compute_competence_cost(action):
# WRITE THE DOSE PENALTY MATH
return ...
# --- 3. MAIN POTENTIAL FUNCTION ---
def potential_function(state, action, t):
# Iterate through features, calculate scores, return weighted sum
# base_potential = ...
decay_factor = time_decay(t)
return base_potential * decay_factor
# --- 4. REWARD FUNCTION ---
def reward_function(s, a, t, s_next, a_next, t_next, gamma=0.99):
return gamma * potential_function(s_next, a_next, t_next) - potential_function(s, a, t) - compute_competence_cost(a)
### DATASET SUMMARY:
"total_records":
"total_patients":
"mortality_rate":
"features_analyzed":
### CRITICAL STATE FEATURES:
1. **feature1**
- Missing:
- Mean:
- Std:
- Range:
- Median:
- Q25: Q75:
- Correlation with mortality:
**Use these statistics along with clinical knowledge to define target range**
**Target should reflect homeostasis: values in normal/healthy physiological range**
**CRITICAL REQUIREMENTS FOR EXECUTABLE CODE:**
- Include ALL imports needed: import numpy as np‘, import pandas as pd‘, import math (if needed)
- Define ALL helper functions you use (e.g., distance functions, penalty functions, cost functions)
- Define ALL constants/variables used (e.g., FEATURE_WEIGHTS‘, FEATURE_TARGETS‘, TAU‘, ACTION_COST_SCALE‘)
- Code must be directly runnable without any undefined variables or functions
- The reward function MUST call potential_function and compute the difference
- ALL imports, helper functions, and constants MUST be defined in your code
- Code must be directly executable - no undefined variables or functions
Appendix B Reward Examples
B.1 Sample of Reward Function for Sepsis Management
1import numpy as np
2import math
3
4# --- 1. PARAMETER DEFINITIONS (HARD-CODED FROM DATA) ---
5
6# Normalization assumptions (min-max from raw data):
7# sofa_24hours: 0-23
8# baseexcess: -25 to 0
9# lactate: 0.3 to 29
10# urineoutput: -3000 to 4400
11# mbp: 20 to 200
12# heartrate: 23 to 212
13
14# Survival config combines best from champions:
15# - Use bell curves for "Goldilocks" features with targets near clinical medians/normals
16# - Use directional decay for directional features (sofa_24hours, urineoutput)
17# - Use decay_low or decay_high style from Champion 3 and 4 for directional features
18SURVIVAL_CONFIG = {
19 ’sofa_24hours’: {
20 ’type’: ’directional_decay’,
21 ’direction’: ’low’, # lower sofa better
22 ’min’: 0.0,
23 ’max’: 1.0,
24 # decay steepness chosen so score ~0.1 at val=1 (max sofa)
25 ’k’: 2.3
26 },
27 ’baseexcess’: {
28 ’type’: ’bell’,
29 ’target’: normalize_raw(-2.0, -25.0, 0.0), # ~0.92
30 ’sigma’: 0.1 # moderate spread from Champion 1 and 4
31 },
32 ’lactate’: {
33 ’type’: ’decay_lower’, # penalize values above target
34 ’target’: normalize_raw(1.6, 0.3, 29.0), # ~0.045
35 ’sigma’: 0.05 # for decay rate calculation
36 },
37 ’urineoutput’: {
38 ’type’: ’directional_decay’,
39 ’direction’: ’high’, # higher urine output better
40 ’threshold’: normalize_raw(40.0, -3000.0, 4400.0), # ~0.414
41 ’k’: 5.0 # steep decay below threshold
42 },
43 ’mbp’: {
44 ’type’: ’bell’,
45 ’target’: normalize_raw(75.0, 20.0, 200.0), # ~0.31
46 ’sigma’: 0.1
47 },
48 ’heartrate’: {
49 ’type’: ’bell’,
50 ’target’: normalize_raw(85.0, 23.0, 212.0), # ~0.33
51 ’sigma’: 0.1
52 },
53}
54
55# Confidence decay taus (hours) from Champion 2 (best confidence anchor)
56# Uniform tau = 6 hours for all features to reflect faster confidence decay
57CONFIDENCE_TAU = {
58 ’sofa_24hours’: 6.0,
59 ’baseexcess’: 6.0,
60 ’lactate’: 6.0,
61 ’urineoutput’: 6.0,
62 ’mbp’: 6.0,
63 ’heartrate’: 6.0,
64}
65
66# Action penalty parameters from Champion 3 (best competence anchor)
67MAX_DOSE_LEVEL = 4
68ACTION_COST_SCALE = 0.25 # penalty scale per action component
69
70# Weights for components balanced as in Knee Point Champion (equal weighting)
71COMPONENT_WEIGHTS = {
72 ’survival’: 1/3,
73 ’confidence’: 1/3,
74 ’competence’: 1/3,
75}
76
77# --- 2. MATH HELPER FUNCTIONS (NO PLACEHOLDERS) ---
78
79def time_decay(t):
80 """
81 Strategic annealing time decay function.
82 Exponential decay with half-life 48 hours.
83 """
84 half_life = 48.0
85 decay = 0.5 ** (t / half_life)
86 return decay
87
88def compute_survival_score(val, params):
89 """
90 Compute survival score for a single normalized feature value in [0,1].
91 Types:
92 - ’bell’: Gaussian bell curve centered at target with sigma.
93 - ’decay_lower’: exponential decay if val > target.
94 - ’directional_decay’: exponential decay from threshold or zero depending on direction.
95 Returns score in [0,1].
96 """
97 if val is None:
98 # Missing value: neutral survival score 0.5
99 return 0.5
100
101 ftype = params[’type’]
102
103 if ftype == ’bell’:
104 target = params[’target’]
105 sigma = params[’sigma’]
106 if sigma <= 0:
107 return 0.0
108 diff = val - target
109 score = math.exp(-0.5 * (diff / sigma) ** 2)
110 return score
111
112 elif ftype == ’decay_lower’:
113 # Penalize values above target exponentially
114 target = params[’target’]
115 sigma = params[’sigma’]
116 if val <= target:
117 return 1.0
118 else:
119 # decay rate so score ~0.5 at val=target+sigma
120 decay_rate = math.log(2) / sigma
121 score = math.exp(-decay_rate * (val - target))
122 return score
123
124 elif ftype == ’directional_decay’:
125 direction = params.get(’direction’, None)
126 if direction == ’low’:
127 # best at 0, decay as val increases
128 k = params.get(’k’, 2.3)
129 score = math.exp(-k * val)
130 return score
131 elif direction == ’high’:
132 threshold = params.get(’threshold’, 0.5)
133 k = params.get(’k’, 5.0)
134 if val >= threshold:
135 return 1.0
136 else:
137 diff = (threshold - val) / threshold if threshold > 0 else 1.0
138 score = math.exp(-k * diff)
139 return score
140 else:
141 # Unknown direction, neutral
142 return 0.5
143 else:
144 # Unknown type, neutral
145 return 0.5
146
147def compute_confidence_weight(delta_t, tau):
148 """
149 Exponential decay of confidence with delta_t (hours).
150 Returns weight in (0,1].
151 """
152 if delta_t is None or delta_t < 0:
153 delta_t = 0
154 return math.exp(-delta_t / tau)
155
156def compute_competence_cost(action):
157 """
158 Compute competence penalty from actions.
159 Penalize higher doses linearly scaled.
160 Sum penalties for vaso and iv fluid, capped at 1.
161 Returns penalty in [0,1].
162 """
163 vaso_level = action.get(’vaso_5quantile’, 0)
164 iv_level = action.get(’iv_fluid_5quantile’, 0)
165 vaso_norm = vaso_level / MAX_DOSE_LEVEL
166 iv_norm = iv_level / MAX_DOSE_LEVEL
167 total_penalty = vaso_norm * ACTION_COST_SCALE + iv_norm * ACTION_COST_SCALE
168 total_penalty = min(total_penalty, 1.0)
169 return total_penalty
170
171# --- 3. MAIN POTENTIAL FUNCTION ---
172
173def potential_function(state, t):
174 """
175 Compute potential function Phi(s,a,t) as balanced weighted sum of survival,
176 confidence, and competence components with strategic time decay.
177 state: dict of feature: (value_normalized, delta_t)
178 action: dict with keys [’vaso_5quantile’, ’iv_fluid_5quantile’]
179 t: absolute time step (int)
180 """
181 survival_scores = []
182 confidence_weights = []
183
184 for feat in SURVIVAL_CONFIG.keys():
185 val, delta_t = state.get(feat, (None, None))
186 surv_score = compute_survival_score(val, SURVIVAL_CONFIG[feat])
187 tau = CONFIDENCE_TAU.get(feat, 6.0)
188 conf_weight = compute_confidence_weight(delta_t, tau) if delta_t is not None else 0.0
189
190 survival_scores.append(surv_score * conf_weight)
191 confidence_weights.append(conf_weight)
192
193 # Aggregate survival component: weighted average survival weighted by confidence
194 sum_confidence = sum(confidence_weights)
195 if sum_confidence > 0:
196 survival_component = sum(survival_scores) / sum_confidence
197 else:
198 survival_component = 0.5 # neutral if no confidence
199
200 # Confidence component: average confidence weight normalized [0,1]
201 if len(confidence_weights) > 0:
202 confidence_component = sum(confidence_weights) / len(confidence_weights)
203 else:
204 confidence_component = 0.0
205
206 # Combine components equally weighted (balanced)
207 base_potential = (
208 COMPONENT_WEIGHTS[’survival’] * survival_component +
209 COMPONENT_WEIGHTS[’confidence’] * confidence_component +
210 COMPONENT_WEIGHTS[’competence’] * competence_component
211 )
212
213 # Clamp base potential to [0,1]
214 base_potential = max(0.0, min(1.0, base_potential))
215
216 # Apply strategic time decay
217 decay_factor = time_decay(t)
218
219 return base_potential * decay_factor
220
221# --- 4. REWARD FUNCTION ---
222
223def reward_function(s, t, s_next, t_next, a, gamma=0.99):
224 """
225 Reward is difference-based with discount factor:
226 R(s,t,s’,t’) = gamma * Phi(s’,t’) - Phi(s,t)
227 """
228 return gamma * potential_function(s_next, t_next) - potential_function(s, t) - compute_competence_cost(a)
B.2 Sample of Reward Function for Ventilation Setting
1import numpy as np
2import math
3
4# --- 1. PARAMETER DEFINITIONS (HARD-CODED FROM DATA) ---
5
6# Feature names: [’mbp’, ’sofa_24hours’, ’lactate’, ’RASS’, ’pH’, ’spo2’, ’paco2’]
7# All features normalized 0-1, so targets and sigmas are also normalized accordingly.
8
9# We define targets and sigmas for bell curves ("Goldilocks" features)
10# and directional decay parameters for directional features.
11
12# Normal physiological ranges and clinical knowledge:
13# mbp (mean blood pressure): normal ~ 65-105 mmHg
14# sofa_24hours: lower is better (0 best, 24 worst)
15# lactate: normal < 2 mmol/L (lower better)
16# RASS: target around -2 (sedation level)
17# pH: normal 7.35-7.45 (center ~7.4)
18# spo2: higher better, normal > 95%
19# paco2: normal 35-45 mmHg (lower better)
20
21# Since features are normalized 0-1, we must map clinical targets to normalized scale.
22# We assume min-max normalization per feature as per dataset range.
23
24# Raw ranges from dataset (for normalization):
25RANGES = {
26 ’mbp’: (1.0, 250.0),
27 ’sofa_24hours’: (0.0, 23.0),
28 ’lactate’: (0.2, 30.0),
29 ’RASS’: (-5.0, 4.0),
30 ’pH’: (6.59, 7.78),
31 ’spo2’: (0.0, 100.0),
32 ’paco2’: (16.9, 157.8),
33}
34
35# Targets and sigmas for bell curve features (normalized)
36# We choose sigma to cover roughly the interquartile range normalized.
37
38# mbp: target ~ 80 mmHg (normal 65-105), normalized:
39mbp_target = normalize(80, *RANGES[’mbp’])
40mbp_sigma = (normalize(105, *RANGES[’mbp’]) - normalize(65, *RANGES[’mbp’])) / 2 # half IQR approx
41
42# sofa_24hours: lower better, target near 0, directional decay
43# lactate: lower better, directional decay
44# RASS: target -2, normalized:
45rass_target = normalize(-2, *RANGES[’RASS’])
46rass_sigma = 0.05 # tight around -2
47
48# pH: target 7.4, normalized:
49ph_target = normalize(7.4, *RANGES[’pH’])
50ph_sigma = (normalize(7.45, *RANGES[’pH’]) - normalize(7.35, *RANGES[’pH’])) / 2
51
52# spo2: higher better, directional decay
53# paco2: lower better, directional decay
54
55# Define survival config with type and parameters:
56# ’bell’ for Goldilocks (bell curve)
57# ’decay_low’ for directional decay where lower is better
58# ’decay_high’ for directional decay where higher is better
59
60SURVIVAL_CONFIG = {
61 ’mbp’: {
62 ’type’: ’bell’,
63 ’target’: mbp_target,
64 ’sigma’: mbp_sigma,
65 ’weight’: 1.0,
66 },
67 ’sofa_24hours’: {
68 ’type’: ’decay_low’, # lower better
69 ’tau’: 0.3,
70 ’weight’: 1.0,
71 },
72 ’lactate’: {
73 ’type’: ’decay_low’, # lower better
74 ’tau’: 0.2,
75 ’weight’: 1.0,
76 },
77 ’RASS’: {
78 ’type’: ’bell’,
79 ’target’: rass_target,
80 ’sigma’: rass_sigma,
81 ’weight’: 0.8,
82 },
83 ’pH’: {
84 ’type’: ’bell’,
85 ’target’: ph_target,
86 ’sigma’: ph_sigma,
87 ’weight’: 1.0,
88 },
89 ’spo2’: {
90 ’type’: ’decay_high’, # higher better
91 ’tau’: 0.3,
92 ’weight’: 1.0,
93 },
94 ’paco2’: {
95 ’type’: ’decay_low’, # lower better
96 ’tau’: 0.25,
97 ’weight’: 1.0,
98 },
99}
100
101# Confidence decay taus for delta_t (hours)
102# Trust decays exponentially with time gap since last real record
103CONFIDENCE_TAU = {
104 ’mbp’: 12,
105 ’sofa_24hours’: 24,
106 ’lactate’: 8,
107 ’RASS’: 6,
108 ’pH’: 12,
109 ’spo2’: 6,
110 ’paco2’: 8,
111}
112
113# Action cost scale: penalty per action level (higher level = higher penalty)
114# Actions: [’PEEP_level’, ’FiO2_level’, ’Tidal_level’]
115# Assume max levels 0-3 (4 levels), scale penalty linearly 0 to 1
116ACTION_MAX_LEVEL = {
117 ’PEEP_level’: 3,
118 ’FiO2_level’: 3,
119 ’Tidal_level’: 3,
120}
121ACTION_COST_WEIGHT = 0.5 # weight of competence penalty in total potential
122
123# --- 2. MATH HELPER FUNCTIONS (NO PLACEHOLDERS) ---
124
125def time_decay(t):
126 """
127 Strategic annealing decay: potential decays toward zero as episode progresses.
128 Use exponential decay with half-life of 48 hours (2 days).
129 """
130 half_life = 48.0
131 decay = 0.5 ** (t / half_life)
132 return decay
133
134def compute_survival_score(val, params):
135 """
136 Compute survival score for one feature normalized val in [0,1].
137 params: dict with keys depending on type:
138 - ’type’: ’bell’, ’decay_low’, ’decay_high’
139 - For bell: ’target’, ’sigma’
140 - For decay: ’tau’
141 Returns score in [0,1].
142 """
143 if params[’type’] == ’bell’:
144 # Gaussian bell curve centered at target with sigma
145 diff = val - params[’target’]
146 score = math.exp(-0.5 * (diff / params[’sigma’])**2)
147 return score
148 elif params[’type’] == ’decay_low’:
149 # Lower values better, exponential decay from 0
150 # score = exp(-val / tau)
151 score = math.exp(-val / params[’tau’])
152 return score
153 elif params[’type’] == ’decay_high’:
154 # Higher values better, exponential decay from 1
155 # score = exp(-(1 - val) / tau)
156 score = math.exp(-(1 - val) / params[’tau’])
157 return score
158 else:
159 # Unknown type, return neutral 0.5
160 return 0.5
161
162def compute_confidence_weight(delta_t, tau):
163 """
164 Exponential decay of confidence with delta_t (hours).
165 """
166 return math.exp(-delta_t / tau)
167
168def compute_competence_cost(action):
169 """
170 Compute penalty cost for actions.
171 Higher dose (level) => higher penalty.
172 Normalize each action level by max level, sum and scale.
173 Return cost in [0,1].
174 """
175 total_cost = 0.0
176 n_actions = len(action)
177 for k, v in action.items():
178 max_level = ACTION_MAX_LEVEL.get(k, 3)
179 norm_level = v / max_level if max_level > 0 else 0
180 total_cost += norm_level
181 avg_cost = total_cost / n_actions if n_actions > 0 else 0
182 # Scale by weight
183 return avg_cost * ACTION_COST_WEIGHT
184
185# --- 3. MAIN POTENTIAL FUNCTION ---
186
187def potential_function(state, t):
188 """
189 Compute potential function Phi(s,t) = survival * confidence
190 with strategic time decay.
191 state: dict of feature -> (val, delta_t)
192 action: dict of action_name -> level
193 t: absolute time step (int)
194 """
195 survival_scores = []
196 confidence_weights = []
197 weights = []
198 for feat, (val, delta_t) in state.items():
199 if feat not in SURVIVAL_CONFIG or feat not in CONFIDENCE_TAU:
200 continue
201 params = SURVIVAL_CONFIG[feat]
202 tau_conf = CONFIDENCE_TAU[feat]
203 survival = compute_survival_score(val, params)
204 confidence = compute_confidence_weight(delta_t, tau_conf)
205 w = params[’weight’]
206 survival_scores.append(survival * confidence * w)
207 weights.append(w)
208 # Normalize survival component by sum weights to balance
209 if weights:
210 survival_component = sum(survival_scores) / sum(weights)
211 else:
212 survival_component = 0.0
213
214 base_potential = survival_component
215 # Clip potential to [0,1] to avoid negative or >1 values
216 base_potential = max(0.0, min(1.0, base_potential))
217
218 decay_factor = time_decay(t)
219
220 return base_potential * decay_factor
221
222# --- 4. REWARD FUNCTION ---
223
224def reward_function(s, t, s_next, t_next, a, gamma=0.99):
225 """
226 Reward = gamma * Phi(s’,t’) - Phi(s,t)
227 """
228 return gamma * potential_function(s_next, t_next) - potential_function(s, t) - compute_competence_cost(a)
B.3 Sample of Reward Function for Renal Replacement Therapy
1import numpy as np
2import math
3
4# --- 1. PARAMETER DEFINITIONS (HARD-CODED FROM DATA) ---
5
6# Feature targets and healthy ranges (normalized 0-1)
7# We define "Goldilocks" features with bell curves centered at median (healthy target)
8# Directional features with monotonic decay from healthy boundary
9
10# From data summary and clinical knowledge:
11# pH normal range ~7.35-7.45 (normalized approx 0.45-0.55)
12# potassium normal ~3.5-5.0 mEq/L (normalized approx 0.25-0.5)
13# mbp (mean blood pressure) normal ~65-105 mmHg (normalized approx 0.3-0.6)
14# creatinine normal low (kidney function) ~0.6-1.2 mg/dL (normalized approx 0.1-0.2)
15# urineoutput normal > 0.5 ml/kg/hr (normalized approx >0.04) but here median 80/2000 normalized ~0.04
16# sofa_24hours lower is better (0 is best), so directional decay from 0 upwards
17# heartrate normal ~60-100 bpm (normalized approx 0.3-0.5)
18
19# We use normalized values 0-1 as input, so we set targets accordingly:
20# For bell curve features: center=median normalized, sigma chosen to cover IQR roughly
21# For directional features: exponential decay from healthy boundary
22
23# Weights to balance components (sum to 1)
24SURVIVAL_WEIGHTS = {
25 ’pH’: 1.0,
26 ’potassium’: 1.0,
27 ’mbp’: 1.0,
28 ’creatinine’: 1.0,
29 ’urineoutput’: 1.0,
30 ’sofa_24hours’: 1.0,
31 ’heartrate’: 1.0,
32}
33# Normalize weights so sum = 1
34total_w = sum(SURVIVAL_WEIGHTS.values())
35for k in SURVIVAL_WEIGHTS:
36 SURVIVAL_WEIGHTS[k] /= total_w
37
38# Targets and sigmas for bell curve (Goldilocks) features
39# Using median and IQR from data summary, normalized 0-1 scale assumed input
40# We approximate sigma as (Q75-Q25)/1.35 (approx std dev for normal dist)
41SURVIVAL_CONFIG = {
42 # bell curve features: pH, potassium, mbp, creatinine, heartrate
43 ’pH’: {
44 ’type’: ’bell’,
45 ’target’: 0.52, # approx normalized median 7.39 in range 6.72-7.94
46 ’sigma’: 0.03, # approx (0.55-0.45)/1.35 ~0.07, but tighter for pH
47 },
48 ’potassium’: {
49 ’type’: ’bell’,
50 ’target’: 0.42, # median 4.2 in range 2.5-9.03 normalized ~0.42
51 ’sigma’: 0.07, # (0.46-0.39)/1.35 ~0.05, slightly relaxed
52 },
53 ’mbp’: {
54 ’type’: ’bell’,
55 ’target’: 0.40, # median 79.67 in range 20.67-198 normalized ~0.4
56 ’sigma’: 0.07, # (0.45-0.36)/1.35 ~0.07
57 },
58 ’creatinine’: {
59 ’type’: ’bell’,
60 ’target’: 0.21, # median 0.211 in range 0.024-1.97 normalized ~0.21
61 ’sigma’: 0.05, # (0.28-0.16)/1.35 ~0.09, tighter penalty for kidney
62 },
63 ’heartrate’: {
64 ’type’: ’bell’,
65 ’target’: 0.43, # median 91.13 in range 0-207 normalized ~0.44
66 ’sigma’: 0.1, # (0.5-0.38)/1.35 ~0.09
67 },
68 # directional decay features:
69 # urineoutput: higher is better, so decay if below threshold
70 ’urineoutput’: {
71 ’type’: ’directional’,
72 ’threshold’: 0.04, # normalized median ~80/1975 ~0.04
73 ’direction’: ’above’, # potential decays if below threshold
74 ’tau’: 0.1,
75 },
76 # sofa_24hours: lower is better, decay if above threshold
77 ’sofa_24hours’: {
78 ’type’: ’directional’,
79 ’threshold’: 0.05, # normalized median 0/14 ~0, so threshold small
80 ’direction’: ’below’, # potential decays if above threshold
81 ’tau’: 10.0,
82 },
83}
84
85# Confidence decay taus for delta_t (hours)
86# Trust decays exponentially with time gap since last real record
87# Use feature-specific taus reflecting clinical urgency
88CONFIDENCE_TAU = {
89 ’pH’: 12, # pH changes fast, trust decays quickly
90 ’potassium’: 24, # potassium changes slower
91 ’mbp’: 6, # blood pressure changes fast
92 ’creatinine’: 48, # kidney function slower
93 ’urineoutput’: 24,
94 ’sofa_24hours’: 72,
95 ’heartrate’: 6,
96}
97
98# Competence penalty scale for action dose levels
99# action ’action’ levels: 0 (no dose) to 3 (high dose)
100# Penalty increases with dose level
101ACTION_COST_SCALE = 0.1 # max penalty 0.3 for dose=3
102
103# --- 2. MATH HELPER FUNCTIONS (NO PLACEHOLDERS) ---
104
105def time_decay(t):
106 """
107 Strategic annealing decay: potential decays toward zero as episode progresses.
108 Use exponential decay with half-life of 100 time steps.
109 """
110 half_life = 100
111 decay = 0.5 ** (t / half_life)
112 return decay
113
114def compute_survival_score(val, params):
115 """
116 Compute survival score for one feature value normalized 0-1.
117 For bell curve: Gaussian centered at target with sigma.
118 For directional: exponential decay from threshold.
119 Returns score in [0,1].
120 """
121 if params[’type’] == ’bell’:
122 target = params[’target’]
123 sigma = params[’sigma’]
124 # Gaussian bell curve
125 diff = val - target
126 score = math.exp(-0.5 * (diff / sigma) ** 2)
127 return score
128 elif params[’type’] == ’directional’:
129 threshold = params[’threshold’]
130 tau = params[’tau’]
131 direction = params[’direction’]
132 if direction == ’above’:
133 # score = 1 if val >= threshold, else decays exponentially below threshold
134 if val >= threshold:
135 return 1.0
136 else:
137 return math.exp(-(threshold - val) / tau)
138 elif direction == ’below’:
139 # score = 1 if val <= threshold, else decays exponentially above threshold
140 if val <= threshold:
141 return 1.0
142 else:
143 return math.exp(-(val - threshold) / tau)
144 else:
145 # Unknown direction, return 0
146 return 0.0
147 else:
148 return 0.0
149
150def compute_confidence_weight(delta_t, tau):
151 """
152 Exponential decay of confidence with delta_t (hours).
153 """
154 return math.exp(-delta_t / tau)
155
156def compute_competence_cost(action):
157 """
158 Penalize higher dose actions.
159 action is dict with key ’action’ and integer dose level 0-3.
160 Returns penalty in [0, 1].
161 """
162 dose = action.get(’action’, 0)
163 # Clamp dose to 0-3
164 dose = max(0, min(3, dose))
165 penalty = dose * ACTION_COST_SCALE
166 # Normalize penalty to max 1 if dose=10 (not expected here)
167 return min(penalty, 1.0)
168
169# --- 3. MAIN POTENTIAL FUNCTION ---
170
171def potential_function(state, t):
172 """
173 Compute potential function Phi(s,t) as weighted sum of survival * confidence - competence cost,
174 then apply strategic time decay.
175 """
176 survival_sum = 0.0
177 confidence_sum = 0.0
178 total_weight = 0.0
179
180 for feat, weight in SURVIVAL_WEIGHTS.items():
181 if feat not in state:
182 # Missing feature, skip
183 continue
184 val, delta_t = state[feat]
185 params = SURVIVAL_CONFIG.get(feat)
186 if params is None:
187 continue
188 survival_score = compute_survival_score(val, params)
189 tau = CONFIDENCE_TAU.get(feat, 24) # default tau=24 if missing
190 confidence_weight = compute_confidence_weight(delta_t, tau)
191 # Combine survival and confidence multiplicatively
192 component_score = survival_score * confidence_weight
193 survival_sum += component_score * weight
194 total_weight += weight
195
196 # Normalize survival component (should be close to 1)
197 if total_weight > 0:
198 survival_component = survival_sum / total_weight
199 else:
200 survival_component = 0.0
201
202 # Apply strategic time decay
203 decay_factor = time_decay(t)
204
205 potential = base_potential * decay_factor
206
207 # Clamp potential to [0,1]
208 potential = max(0.0, min(1.0, potential))
209
210 return potential
211
212# --- 4. REWARD FUNCTION ---
213
214def reward_function(s, t, s_next, t_next, a, gamma=0.99):
215 """
216 Reward is difference-based with discount factor:
217 R = gamma * Phi(s’,t’) - Phi(s,t)
218 """
219 phi_next = potential_function(s_next, t_next)
220 phi_curr = potential_function(s, t)
221 reward = gamma * phi_next - phi_curr - compute_competence_cost(a)
222 return reward
Appendix C Mathematical form of Drives in Reward Function
C.1 Survival Score $\mathcal{S}$
Three canonical forms for normalized feature values:
[…]
  • Directional Decay: $S_{\text{directional}}(x)=\begin{cases}\exp\left(-\dfrac{x}{\tau}\right)&\text{(lower-better)}\\ \exp\left(-\dfrac{1-x}{\tau}\right)&\text{(higher-better)}\end{cases}$
  • Asymmetric Decay-Lower (thresholded): $S_{\text{asymetric}}(x)=\begin{cases}1&x\leq\mu\\ \exp\left(-\dfrac{\ln 2}{\sigma}(x-\mu)\right)&x>\mu\end{cases}$
C.2 Confidence Weight $\mathcal{U}$
$U(\Delta t;\tau)=\exp\left(-\frac{\Delta t}{\tau}\right)$
C.3 Competence Cost $\mathcal{C}$
$C(\mathbf{a})=w_{c}\sum_{i=1}^{N}\frac{a_{i}}{a_{\max,i}}$
Appendix D Theoretical Proof
D.1 Breaking Policy Invariance
Proposition D.1 (Cost Regularization Breaks Invariance and Enforces Efficiency).
Let the reward function be defined as $R(s_{t},a_{t},s_{t+1})=F(s_{t},s_{t+1})-\lambda\mathcal{C}(a_{t})$, where $F(s_{t},s_{t+1})=\gamma\Phi(s_{t+1})-\Phi(s_{t})$ is a potential-based shaping function and $\mathcal{C}(a_{t})>0$ is a strictly positive action cost.
While the shaping term $F$ forms a telescoping sum that preserves policy invariance (Ng et al., 1999), the inclusion of the non-telescoping cost term $\lambda\mathcal{C}(a_{t})$ renders the cumulative return path-dependent. This modification compels the optimal policy to satisfy two objectives: maximizing the time-discounted physiological recovery $\gamma^{T}\Phi(s_{T})$ (preventing prolonged ICU stays) while simultaneously minimizing the cumulative intervention cost.
Proof.
Consider a trajectory $\tau=(s_{0},a_{0},s_{1},\dots,s_{T})$. The discounted cumulative return $G(\tau)$ is:
$G(\tau)=\sum_{t=0}^{T-1}\gamma^{t}\left[\left(\gamma\Phi(s_{t+1})-\Phi(s_{t})\right)-\lambda\mathcal{C}(a_{t})\right].$ (11)
We separate the potential-based shaping term from the cost term. The shaping term telescopes as follows:
$\sum_{t=0}^{T-1}\gamma^{t}(\gamma\Phi(s_{t+1})-\Phi(s_{t}))$ $=\sum_{t=0}^{T-1}(\gamma^{t+1}\Phi(s_{t+1})-\gamma^{t}\Phi(s_{t}))$ (12)
$=\gamma^{T}\Phi(s_{T})-\Phi(s_{0}).$ (13)
Substituting this back into $G(\tau)$, we obtain the decomposed objective:
$G(\tau)=\underbrace{\left(\gamma^{T}\Phi(s_{T})-\Phi(s_{0})\right)}_{\text{Telescoped Potential (Time-Dependent)}}-\underbrace{\lambda\sum_{t=0}^{T-1}\gamma^{t}\mathcal{C}(a_{t})}_{\text{Cumulative Cost (Path-Dependent)}}.$ (14)
Here, the first term depends only on the boundary states and the trajectory duration $T$. The factor $\gamma^{T}$ provides a strategic incentive to reduce $T$ (avoiding ”holding” behavior), but remains invariant to the path taken to reach $s_{T}$ at time $T$.
However, the second term, $\lambda\sum\gamma^{t}\mathcal{C}(a_{t})$, strictly does not telescope. It introduces a dependency on the specific sequence of actions $a_{0:T-1}$. Consequently, among policies that achieve the same recovery state $\Phi(s_{T})$ in the same time $T$, the objective function breaks invariance to strictly prefer the policy with the minimum cumulative intervention cost. ∎
D.2 Equivalence to Lagrangian Relaxation of CMDP
Proposition D.2 (Equivalence to Lagrangian Relaxation of CMDP).
Maximizing the expected cumulative return of the regularized reward function $R(s,a,s^{\prime})=(\gamma\Phi(s^{\prime})-\Phi(s))-\lambda\mathcal{C}(a)$ is equivalent to maximizing the Lagrangian relaxation of a Constrained Markov Decision Process (CMDP), where the objective is to maximize physiological gain $J_{\Phi}$ subject to a constraint on the cumulative intervention cost $J_{\mathcal{C}}\leq\beta$.
Proof.
We define the primal CMDP optimization problem as:
Maximize: $J_{\Phi}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\gamma\Phi(s_{t+1})-\Phi(s_{t})\right)\right]$ (15)
Subject to: $J_{\mathcal{C}}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}(a_{t})\right]\leq\beta.$ (16)
The Lagrangian function $\mathcal{L}(\pi,\lambda)$ with Lagrange multiplier $\lambda\geq 0$ is constructed as:
$\mathcal{L}(\pi,\lambda)=J_{\Phi}(\pi)-\lambda(J_{\mathcal{C}}(\pi)-\beta).$ (17)
Expanding the expectations and rearranging terms by the linearity of the summation:
$\mathcal{L}(\pi,\lambda)$ $=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\gamma\Phi(s_{t+1})-\Phi(s_{t})\right)-\lambda\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}(a_{t})\right]+\lambda\beta$ (18)
$=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\underbrace{\gamma\Phi(s_{t+1})-\Phi(s_{t})-\lambda\mathcal{C}(a_{t})}_{R(s_{t},a_{t},s_{t+1})}\right)\right]+\text{const.}$ (19)
Thus, finding the optimal policy $\pi^{*}$ that maximizes the Lagrangian $\mathcal{L}(\pi,\lambda)$ for a fixed $\lambda$ corresponds exactly to maximizing the value function defined by the composite reward $R$. The parameter $\lambda$ serves as the dual variable governing the trade-off between physiological stabilization and intervention cost. ∎
Appendix E Hyperparameters and Experimental Details
We separate the data into 7:1:1:1 split on patient level as training set for policy, training set for reward, test set for policy and test set for reward. All the experiments are run with 5 different seeds.
E.1 Computation Resources
All experiments were conducted on a high-performance computing cluster node equipped with 8 $\times$ NVIDIA H200 NVL GPUs.
E.2 Selection of LLM
For the CodeGen baselines, we intentionally selected the most powerful state-of-the-art proprietary models available. Since the baseline approach involves a simple, single-pass inference to generate a policy, the computational cost is manageable even with expensive APIs. This ensures our method is compared against the strongest possible ”zero-shot” clinical agents, establishing a rigorous upper bound for performance without regarding training constraints. In contrast, our medR framework requires generating dense reward signals across vast state-action spaces, making the use of commercial APIs prohibitively expensive and computationally intractable for large-scale training. Consequently, we employ smaller, open-source models that can be deployed locally. A key finding of our work is that medR, even when powered by these smaller, accessible models, consistently outperforms baselines using significantly larger, proprietary LLMs. This validates that our improvements stem from the theoretical rigor of our potential-based Lagrangian framework rather than raw model scale.
E.3 Reward Function Hyperparameters
The discount factor $\gamma$ was set to 0.99 for all environments. For feature selection we use $\kappa=0.6$ for the threshold and the number of features selected $k=7$. We generate $N=20$ functions as reward candidates for selection. For $J_{surv}$ epsilon is 2. For $J_{comp}$ we use $k=10$ for $H$, $\alpha=0.1$ for $E$.
E.4 RL Policy Hyperparameters
Sepsis Treatment
The Sepsis task utilizes the MIMIC-III dataset with a state dimension of 46 and a discrete action space of 25. The policy network is constructed with 2 hidden layers of 32 units each. Optimization was performed with a batch size of 512 and a learning rate of $1\times 10^{-3}$ over 10 epochs.
Mechanical Ventilation
For the Ventilation task, data is sourced from the eICU database. The environment consists of a 41-dimensional state space and a discrete action space of size 18. The network architecture uses 3 hidden layers with 32 units per layer. Training used a batch size of 512, a learning rate of $5\times 10^{-4}$, and ran for 10 epochs.
Renal Replacement Therapy
The RRT task is based on the AmsterdamUMCdb dataset, featuring a compact state dimension of 18 and a binary action space (size 1). The network is deeper but narrower, using 3 hidden layers with 16 units each. Due to the smaller dataset size, a smaller batch size of 256 was used, with a learning rate of $1\times 10^{-3}$ for 10 epochs.
E.5 Feature Selection
For each task, a subset of critical features was identified to inform the LLM-generated reward potential. These features are listed below:
  • Sepsis: SOFA score, Base Excess, Lactate, Urine Output, Mean Blood Pressure, Heart Rate.
  • Ventilation: Mean Blood Pressure, SOFA score, Lactate, RASS, pH, SpO2, PaCO2.
  • RRT: pH, Potassium, Mean Blood Pressure, Creatinine, Urine Output, SOFA score, Heart Rate.
Appendix F Related Work
F.1 Reward Design in Healthcare RL:
Traditional RL methodologies in healthcare typically define reward functions using clinical indicators, such as changes in SOFA scores, expert-derived heuristics, or terminal outcomes, including 90-day mortality. A foundational example is the AI Clinician by (Komorowski et al., 2018), which employed sparse, outcome-based rewards reflecting 90-day survival. A subsequent analysis(Huang et al., 2022) introduced safety constraints by limiting actions to those frequently taken by clinicians, further restricting exploration. (Raghu et al., 2017) addressed the limitations of sparse signals by incorporating intermediate clinical metrics such as SOFA, lactate, and increasing feedback frequency through their heuristic-based approach, which risks encoding clinical bias. (Raghu et al., 2018) and (Peng et al., 2018) advanced this direction by using model-based shaping, where predicted mortality log-odds serve as reward signals. While this enhances continuity, it introduces dependence on opaque, learned predictors. (Futoma et al., 2017) employed high-fidelity simulators to define reward functions, though such strategies are resource-intensive and may lack clinical transparency. Despite these advances, most reward functions rely on proxy measures that lack semantic clarity and deep clinical grounding. Reward design thus remains an open challenge, necessitating signals that are temporally dense, interpretable, and ethically aligned.
F.2 LLMs in Clinical Settings
LLMs have transformed natural language tasks in clinical medicine (Thirunavukarasu et al., 2023; Haltaufderheide and Ranisch, 2024; Shool et al., 2025), aiding in medical question answering (Singhal et al., 2025), policy critique (Yang et al., 2025), and patient summaries(Asgari et al., 2025). However, their application in structured reward modelling for healthcare RL remains largely unexplored, presenting a promising opportunity. Recent studies have started exploring how LLMs can enhance RL in healthcare by improving sample efficiency and reward design.
For example,(Bhambri et al., 2024) uses LLMs for abstract plans in potential-based reward engineering, boosting sample efficiency in toy environments, but their method relies on handcrafted abstractions and external verifiers.(Wu, 2024) proposes Q-engineering, which utilises LLM heuristics to shape Q-values, thereby improving efficiency but relying on task-specific heuristics and lacking interpretability. (Zhang et al., 2024b) shows how LLMs refine intrinsic rewards through code synthesis in rule-based environments, but their method is constrained by fixed domain rules and static strategies.
These methods advance LLM-guided reward design; however, they remain limited by the use of symbolic models, heuristics, and predefined knowledge, which reduces their applicability in complex clinical settings. These limitations are highlighted by (Gao et al., 2024), who explore reward models but face reward hacking issues.(Pternea et al., 2024) proposes an RL/LLM taxonomy but doesn’t address real-world clinical challenges, while (Sun, 2023) reviews RLHF but struggles with sparse feedback in healthcare. In contrast, our approach leverages LLMs for temporal credit assignment in clinical settings, transforming sparse outcomes into clinically interpretable rewards from raw patient data, without relying on external verifiers or predefined rules, enabling scalable, adaptive learning for dynamic healthcare environments.
Recent studies have made significant strides in LLM-based reward engineering for RL. (Nazir and Banerjee, 2025) use zero-shot LLMs to replace human feedback, improving sample efficiency.(Fu et al., 2025) introduces PAR, addressing reward hacking in RLHF. (Deng et al., 2025) proposes the LMGT framework to balance exploration and exploitation, while (Wang et al., 2025a) develop SPA-RL for stepwise reward attribution. (Zhang et al., 2024b) proposes a simple framework using LLMs for intrinsic reward generation.
Despite these advancements, our approach is unique. Unlike methods that rely on predefined models or biased human feedback, we utilise raw clinical data to generate clinically interpretable rewards, thereby eliminating the need for external verifiers or predefined rules and offering a scalable, adaptive solution for real-time healthcare applications.
F.3 Credit Assignment
Temporal credit assignment remains a major challenge in RL, for complex, long-horizon tasks. Traditional methods, such as TD-lambda ((Sun, 2023)) and GAE ((Schulman et al., 2015)), focus on temporal differences but lack semantic reasoning, which limits their performance. Recent advances, such as Universal GAE by (Kwiatkowski et al., 2023), improve generalisation. Meanwhile, transformer-based TD learning by (Wang et al., 2025b), which integrates attention-based models, addresses some limitations but still falls short in complex, real-world tasks.
LLMs have emerged as powerful solutions, bringing new levels of contextual understanding and reasoning to RL, effectively tackling long-standing challenges in temporal credit assignments. For example, (Pignatelli et al., 2024) introduces CALM, where LLMs act as critics to decompose tasks into subgoals and assign rewards in zero-shot settings; (Jiang et al., 2025) presents QLLM, replacing complex mixing networks in multi-agent RL with LLM-generated, interpretable credit functions; and (Qu et al., 2025) propose LaRe, using LLMs to derive multi-dimensional latent rewards that capture nuanced contributions in episodic tasks. Collectively, these works establish LLMs as scalable, interpretable, and powerful tools for automating credit assignments in both single- and multi-agent environments.
Our approach, medR, uniquely advances this line by leveraging real clinical data to integrate Large Language Models with potential-based reward shaping. Unlike CALM’s reliance on toy subgoals (Pignatelli et al., 2024), QLLM’s multi-agent focus (Pignatelli et al., 2024), or LaRe’s (Qu et al., 2025) latent outcome alignment, medR transforms sparse terminal outcomes into dense, mathematically grounded learning signals. By prompting LLMs to estimate the physiological stability of patient states ($\Phi(s)$), we construct a regularized reward structure that explicitly balances clinical improvement with intervention cost. This formulation solves a Lagrangian-constrained optimization problem, bridging the gap between high-level medical reasoning and precise temporal credit assignment in critical care.
Appendix G Analysis of Tri-Drive Fitness
To assess the reward functions before training, we utilize the three proposed fitness objectives: Survival ($J_{\text{surv}}$), Confidence ($J_{\text{conf}}$), and Competence ($J_{\text{comp}}$). We acknowledge that baseline rewards explicitly designed with sparse outcome signals will naturally exhibit high $J_{\text{surv}}$ scores by definition, as they directly encode the survival target. Consequently, a superior reward function should not merely maximize $J_{\text{surv}}$ but demonstrate a robust balance across all three Pareto dimensions, particularly in Competence ($J_{\text{comp}}$) and Confidence ($J_{\text{conf}}$). We use J in both reward selection and evaluation since it is an off-policy evaluation functional of a fixed dataset and a fixed clinician policy, not a training objective. Reward selection and policy training are decoupled in our setting. We treat reward design as a model selection problem, and J as a task-level performance criterion that is external to the learning dynamics.
Figure 4: Decomposition of Tri-Drive fitness across tasks. The stacked bars represent the contribution of three distinct evaluation metrics: Survival ($J_{surv}$), Confidence ($J_{conf}$), and Competence ($J_{comp}$).
Appendix H Features and Actions
H.1 Sepsis
The sepsis cohort utilizes 46 observation features, comprising demographics, vital signs, and a comprehensive set of laboratory values, as detailed in Table 4. For the action space, we consider two primary interventions: IV fluids and Vasopressors. The dosage for each intervention is discretized into four quantile bins plus a no-treatment option, resulting in 5 discrete levels per drug. The joint action space is the Cartesian product of these levels, yielding $5\times 5=25$ discrete treatment combinations.
Table 4: List of features for the Sepsis task.
Category Feature Name
Demographics (6) Age, Gender, Weight, Readmission, Elixhauser Score, Mechanical Ventilation Status
Vital Signs (8) Heart Rate, Respiratory Rate, SpO2, Temperature, SBP, DBP, MBP, Shock Index
Lab Values (30) Lactate, PaO2, PaCO2, pH, Base Excess, CO2, Hemoglobin, PaO2/FiO2, WBC, Platelet, BUN, Creatinine, PTT, PT, INR, AST, ALT, Bilirubin, Magnesium, Ionized Calcium, Calcium, Urine Output, Potassium, Sodium, Chloride, Glucose, Albumin, Bicarbonate, FiO2
Scores (2) SOFA (24hr), SIRS, GCS
Abbreviations - SBP/DBP/MBP: Systolic/Diastolic/Mean Blood Pressure; GCS: Glasgow Coma Scale; BUN: Blood Urea Nitrogen; WBC: White Blood Cell; PTT: Partial Thromboplastin Time; PT: Prothrombin Time; INR: International Normalized Ratio; AST: Aspartate Aminotransferase; ALT: Alanine Aminotransferase.
H.2 Mechanical Ventilation (MV)
For the ventilation task, we utilize 41 features selected for training, listed in Table 5. The action space consists of 18 discrete combinations derived from three ventilator settings: Positive End-Expiratory Pressure (PEEP), Fraction of Inspired Oxygen (FiO2), and Tidal Volume adjusted for ideal body weight. These are categorized into Low, Medium, and High levels as summarized in Table 6.
Table 5: List of features for the Mechanical Ventilation task.
Category Feature Name
Demographics (4) Age, Gender, Weight, Elixhauser Score
Vital Signs (7) Heart Rate, Respiratory Rate, SpO2, Temperature, SBP, DBP, MBP
Lab Values (26) Lactate, Bicarbonate, PaCO2, pH, Base Excess, Chloride, Potassium, Sodium, Glucose, Hemoglobin, Ionized Calcium, Calcium, Magnesium, Albumin, BUN, Creatinine, WBC, Platelet, PTT, PT, INR, ETCO2, Urine Output, Rate Std, Analgesic/Sedative Admin, Neuromuscular Blocker Admin
Scores (4) GCS, SOFA (24hr), SIRS (24hr), RASS
Table 6: Categorization of action levels for PEEP, FiO2 and Tidal Volume in the MV task.
Intervention Category Threshold
PEEP (cmH2O) Low $\leq 5$
High >5>5
FiO2 (%) Low <35<35
Medium 35355050
High $\geq 50$
Tidal Volume (ml/kg IBW) Low <6.5<6.5
Medium 6.56.588
High $\geq 8$
H.3 Renal Replacement Therapy (RRT)
The RRT task utilizes a specific subset of 18 features focusing on renal function and hemodynamic stability, as shown in Table 7. Unlike the multi-dimensional discrete actions in Sepsis and Ventilation, the RRT action space is defined as a single dimension representing the aggregate dosage of renal replacement therapy.
Table 7: List of features for the RRT task.
Category Feature Name
Demographics (3) Age, Gender, Weight
Vital Signs (5) Heart Rate, SBP, MBP, DBP, SpO2
Lab Values (8) Urea, Creatinine, pH, Bicarbonate, Potassium, Sodium, PaO2/FiO2, SOFA (24hr)
Urine Output (2) Total Urine Output (6hr), Urine Output Change Rate (6hr)
Appendix I Behavior Policy
To approximate the clinician’s decision-making process and facilitate Off-Policy Evaluation (OPE), we developed a behavior policy parameterized by a Multi-Layer Perceptron (MLP). This policy was trained via supervised learning to imitate observed clinical actions. Across all tasks, models were trained for 100 epochs, with architectures and hyperparameters optimized for each specific domain. For Sepsis, we employed a 3-layer network with 32 hidden units (batch size 256, learning rate $5\times 10^{-4}$). The Ventilation task utilized a compact 2-layer architecture with 8 hidden units (batch size 128, learning rate $1\times 10^{-3}$), while the RRT model consisted of 3 layers with 16 hidden units (batch size 128, learning rate $5\times 10^{-4}$).
Appendix J Pseudocodes
Algorithm 1 Reward Engineering via Tri-Drive Potential Functions
Input: Clinical Dataset $\mathcal{D}$, LLM $\mathcal{M}$, Candidate Features $\mathcal{F}$
Output: Pareto-Optimal Set of Reward Functions $\mathcal{R}^{*}$
// Phase 1: Interpretable Feature Selection
 Calculate statistical metadata $M_{f}$ for all $f\in\mathcal{F}$ (missingness, correlation $\rho$)
$F_{crit}\leftarrow\text{EnsembleSelect}(\mathcal{M}(M_{f},\text{Prompt}_{feat}))$
// Phase 2: Candidate Reward Function Generation
 Initialize Candidate Set $\mathcal{C}\leftarrow\emptyset$
for $i=1$ to NN do
  $r_{i}\leftarrow\text{ExtractCode}(\mathcal{M}(\text{Prompt}_{gen},F_{crit}))$
  Add $r_{i}$ to $\mathcal{C}$
end for
// Phase 3: Multi-Objective Selection
 For all $r\in\mathcal{C}$, compute Offline Fitness Vector:
  $\mathbf{J}(\Phi)=[J_{surv},J_{conf},J_{comp}]$
Finalize: $\mathcal{R}^{*}\leftarrow\text{ParetoFront}(\mathcal{C})$
Appendix K Comparison of Action Distributions between RL and Clinicians
Table 8: Agreement rates between AI policy and Clinician actions across three clinical tasks. Values represent Mean $\pm$ Standard Deviation across seeds. The **Joint** column represents the exact match rate across all action dimensions combined. Bold indicates the highest agreement in each category.
Method Sepsis Agreement (%) Ventilation Agreement (%) RRT (%)
Joint IV Vaso Joint PEEP FiO2 Tidal Dose
ORM $23.99\pm 0.40$ $29.06\pm 0.23$ $75.57\pm 0.89$ $21.28\pm 2.20$ $72.97\pm 0.36$ $51.66\pm 5.10$ $47.28\pm 0.40$ $54.39\pm 9.23$
PRM $25.82\pm 0.55$ $31.62\pm 0.71$ $76.34\pm 0.65$ $22.17\pm 2.72$ $75.29\pm 1.64$ $\mathbf{55.29\pm 5.48}$ $46.06\pm 0.50$ $52.54\pm 11.33$
OPRM $25.50\pm 0.54$ $31.13\pm 0.60$ $76.34\pm 0.81$ $20.81\pm 2.50$ $73.01\pm 0.68$ $54.63\pm 2.65$ $42.88\pm 1.79$ $60.43\pm 3.12$
[…]
Figure 5: Sepsis Policy Agreement. Confusion matrices comparing AI vs. Clinician actions for IV Fluids (left) and Vasopressors (right) across different reward functions.
Figure 6: Ventilation Policy Agreement. Comparison of discrete action choices for PEEP, FiO2, and Tidal Volume. Darker diagonals indicate higher concordance with clinical standards.
Figure 7: RRT Dosing Agreement. Analysis of continuous dosing converted to discrete clinical bins (Low, Mid, High). The proposed method shows stronger alignment in the ’Mid’ (Standard) dosing range.

### v2

@@
 By integrating LLM-driven domain knowledge, our framework automates the design of reward functions for specific diseases while significantly enhancing the performance of the resulting policies.
-Machine Learning, ICML
 1 Introduction
 Critical care medicine is fundamentally defined by high-stakes, sequential decision-making, RL is a natural fit for optimizing DTRs where an agent must make sequential decisions that adapts to a patient’s changing state over time.
-Reward serves as the core of RL that not only guides the agent learning but also defines the optimization problem (Zhu et al., ).
-Small changes in reward function weights can cause profound policy changes(Roggeveen et al., 2024) and bad reward signals affect beyond “win-or-lose” in gaming and can lead to “life-or-death” in healthcare.
-Errors in sensitive fields such as autonomous driving have already been shown to be problematic (Knox et al., 2023).
+Reward serves as the core of RL that not only guides the agent learning but also defines the optimization problem Zhu et al.
+.
+Small changes in reward function weights can cause profound policy changesRoggeveen et al.
+(2024) and bad reward signals affect beyond “win-or-lose" in gaming and can lead to “life-or-death" in healthcare.
+Errors in sensitive fields such as autonomous driving have already been shown to be problematic Knox et al.
+(2023).
 However, reward engineering in clinical RL is less studied and benchmarked.
-This underscores an urgent need to design better rewards that align with clinical goals (Basu et al., 2025).
+This underscores an urgent need to design better rewards that align with clinical goals Basu et al.
+(2025).
 Current rewards in RL for DTRs largely depend on manual and static design.
-Drawing an analogy to recent advances in LLM, we categorize existing clinical reward mechanisms into three distinct paradigms: sparse Outcome Reward Model (ORM), which assigns scalar rewards based on terminal patient outcomes like survival (Komorowski et al., 2018).
-The second is dense Process Reward Model (PRM), which depends on dense feedback based on intermediate physiological state changes (Zhang et al., 2024c).
-To mitigate the sparsity of terminal outcomes, the standard approach augments them with intermediate feedback to form a composite reward ($R_{OPRM}=R_{ORM}+R_{PRM}$) (Raghu et al., 2018; Lu et al., 2024).
+Drawing an analogy to recent advances in LLM, we categorize existing clinical reward mechanisms into three distinct paradigms: sparse Outcome Reward Model (ORM), which assigns scalar rewards based on terminal patient outcomes like survival Komorowski et al.
+(2018).
+The second is dense Process Reward Model (PRM), which depends on dense feedback based on intermediate physiological state changes Zhang et al.
+(2024c).
+To mitigate the sparsity of terminal outcomes, the standard approach augments them with intermediate feedback to form a composite reward ($R_{OPRM}=R_{ORM}+R_{PRM}$) Raghu et al.
+(2018); Lu et al.
+(2024).
 In this paradigm, clinicians manually select disease-relevant features and encode them using simple heuristic logic.
-However, the terminal reward is often weighted disproportionately high, overshadowing the dense shaping signals and the selection of specific parameters and weights remains largely ad-hoc, lacking systematic justification (Lu et al., 2024; Zhang et al., 2024c).
+However, the terminal reward is often weighted disproportionately high, overshadowing the dense shaping signals and the selection of specific parameters and weights remains largely ad-hoc, lacking systematic justification Lu et al.
+(2024); Zhang et al.
+(2024c).
 Despite efforts in designing various reward signals to accelerate policy learning, there are currently no guidelines on designing effective rewards for clinical RL due to several major challenges.
@@
 There is a profound gap between the high-dimensional complexity of patient physiology and the scalar reward value required by RL algorithms.
-This gap leads to misaligned objectives, such as the “Pyrrhic Victory”(Livingston et al., 2019) where an agent optimizes for survival via aggressive interventions or reward hacking.
+This gap leads to misaligned objectives, such as the “Pyrrhic Victory”Livingston et al.
+(2019) where an agent optimizes for survival via aggressive interventions or reward hacking.
 Third, rewards are not verifiable.
@@
 Recent advancements have identified LLMs as a promising solution to bridge the gap between high-level human intent and low-level reinforcement signals.
-While early work utilized LLMs as direct evaluators of agent behavior(Kwon et al., 2023), more recent methods focus on using LLMs to generate executable reward function codes(Ma et al., 2023; Xie et al., 2023).
+While early work utilized LLMs as direct evaluators of agent behaviorKwon et al.
+(2023), more recent methods focus on using LLMs to generate executable reward function codesMa et al.
+(2023); Xie et al.
+(2023).
 We leverage this generative capability to construct a reward function that captures the nuance of DTRs without the prohibitively high cost of manual trial-and-error engineering.
@@
 2.1 Intrinsic Motivation
-In this work, our reward function design follows the evolutionary perspective proposed in (Singh et al., 2009).
+In this work, our reward function design follows the evolutionary perspective proposed in Singh et al.
+(2009).
 They distinguish between reward signals and fitness functions where reward signals are internal scalar values that drive learning within an agent and fitness is the external evaluation that measures an agent’s success.
-For bounded agents in an offline environment, the optimal reward function is not the fitness function(Singh et al., 2010).
+For bounded agents in an offline environment, the optimal reward function is not the fitness functionSingh et al.
+(2010).
 The design of reward functions should optimize intrinsic motivations (IM), unlike previous literature about IM that focuses on incentivizing exploration which is dangerous in our offline setting, in our paper, we form it as a threefold drive: the drive to restore homeostasis(survival), the drive to lower uncertainty(confidence), and the drive to minimize cost(competence).
-In Hull’s Drive Reduction Theory(Hull, 1945), organisms are motivated to maintain physiological equilibrium, we treat homeostasis as a proxy of survival to optimize in the reward function.
-Unlike curiosity-driven frameworks that encourage exploration under uncertainty(Berlyne, 1960; Schmidhuber, 1991), in order to design a reward function specific to our offline environment(Singh et al., 2009), we conversely suppress uncertainty due to the sensitivity in critical care(Djulbegovic, 2011).
+In Hull’s Drive Reduction TheoryHull (1945), organisms are motivated to maintain physiological equilibrium, we treat homeostasis as a proxy of survival to optimize in the reward function.
+Unlike curiosity-driven frameworks that encourage exploration under uncertaintyBerlyne (1960); Schmidhuber (1991), in order to design a reward function specific to our offline environmentSingh et al.
+(2009), we conversely suppress uncertainty due to the sensitivity in critical careDjulbegovic (2011).
 White (1959) furthers the learning process of effective interaction with the environment, which is transferrable as effective treatment under clinical setting.
@@
 Instead, we form it as the gradient of potential to capture the patient state change.
-To bridge the gap between sparse outcome labels and dense clinical supervision, the reward function $R$ is composed of three distinct drives—Survival, Confidence, and Competence leveraging the intrinsic motivation theory (Singh et al., 2009).
+To bridge the gap between sparse outcome labels and dense clinical supervision, the reward function $R$ is composed of three distinct drives—Survival, Confidence, and Competence leveraging the intrinsic motivation theory Singh et al.
+(2009).
 We define the reward function as follows:
-$\small R(s,a,t,s^{\prime},a^{\prime},t^{\prime})=\left[\gamma\Phi(s_{t+1},t+1)-\Phi(s_{t},t)\right]-\lambda\mathcal{C}(a_{t})$ (1)
-$\small\Phi(s_{t},t)=\delta(t)\cdot\sum_{f\in\mathcal{F}}\omega_{f}\cdot\mathcal{S}_{f}(v_{t,f})\cdot\mathcal{U}_{f}(\Delta t_{t,f})$ (2)
+$R(s,a,t,s^{\prime},a^{\prime},t^{\prime})=\left[\gamma\Phi(s_{t+1},t+1)-\Phi(s_{t},t)\right]-\lambda\mathcal{C}(a_{t})$ (1)
+$\Phi(s_{t},t)=\delta(t)\cdot\sum_{f\in\mathcal{F}}\omega_{f}\cdot\mathcal{S}_{f}(v_{t,f})\cdot\mathcal{U}_{f}(\Delta t_{t,f})$ (2)
 where $\gamma$ is the discount factor, $\Phi(\cdot)$ is the health potential, $\delta(t)$ is the strategic decay, $\mathcal{S}_{k}(v_{t,f})$ is the survival score with confidence weight $\mathcal{U}_{f}(\Delta t_{t,f})$ and $\mathcal{C}(\cdot)$ is the competence cost.
@@
 Mathematically, this serves as a regularization term: if two treatments result in the same physiological transition $s_{t}\rightarrow s_{t+1}$, the potential function assigns a higher value to the less invasive action, penalizing over-treatment or flailing behavior.
-By decoupling the accumulating action cost $\mathcal{C}(a)$ from the telescoping potential $\Phi(s,t)$D.1, we effectively formulate the RL objective as the Lagrangian of a Constrained Markov Decision Process (CMDP).
-The agent maximizes the Lagrangian $\mathcal{L}=\text{Physiology}-\lambda\cdot\text{Toxicity}$, ensuring that high-stability states are only pursued if the cost of the intervention path does not exceed the value of the stabilizationD.2.
+By decoupling the accumulating action cost $\mathcal{C}(a)$ from the telescoping potential $\Phi(s,t)$1, we effectively formulate the RL objective as the Lagrangian of a Constrained Markov Decision Process (CMDP).
+The agent maximizes the Lagrangian $\mathcal{L}=\text{Physiology}-\lambda\cdot\text{Toxicity}$, ensuring that high-stability states are only pursued if the cost of the intervention path does not exceed the value of the stabilization2.
 Instead of manual heuristic tuning, the LLM acts as the logic designer.
@@
 We formulate reward design as a Tri-Drive Selection Problem aiming to balance three conflicting objectives based on trajectory-level labeled ground truth: Survival, Confidence, and Competence Fitness.Survival Fitness ($J_{surv}$) validates whether the induced reward signal correlates with actual patient outcomes.
-We define a ground-truth trajectory score $G(\tau)$ that consists of both mortality and Sequential Organ Failure Assessment(SOFA) score(Lambden et al., 2019) increase compared to baseline(time when admitted to ICU).
+We define a ground-truth trajectory score $G(\tau)$ that consists of both mortality and Sequential Organ Failure Assessment(SOFA) scoreLambden et al.
+(2019) increase compared to baseline(time when admitted to ICU).
 The fitness is the correlation $\rho$ between the cumulative reward and $G(\tau)$:
@@
 We define the unified homeostasis score $H(x)\in[0,1]$ for a feature of type $T\in\mathcal{T}$ as:
-$\small h(f)=\begin{cases}\mathbf{1}_{f\in\mathcal{I}}+\mathbf{1}_{f\notin\mathcal{I}}\cdot\sigma\left(k\left[0.5-\frac{d(f,\mathcal{I})}{\text{IQR}}\right]\right)&\text{if }T=N,\\ \sigma\left(k\left[0.5-f\right]\right)&\text{if }T=D_{l},\\ \sigma\left(-k\left[0.5-f\right]\right)&\text{if }T=D_{h},\end{cases}$ (7)
+$h(f)=\begin{cases}\mathbf{1}_{f\in\mathcal{I}}+\mathbf{1}_{f\notin\mathcal{I}}\cdot\sigma\left(k\left[0.5-\frac{d(f,\mathcal{I})}{\text{IQR}}\right]\right)&\text{if }T=N,\\ \sigma\left(k\left[0.5-f\right]\right)&\text{if }T=D_{l},\\ \sigma\left(-k\left[0.5-f\right]\right)&\text{if }T=D_{h},\end{cases}$ (7)
 $H(s_{t})=\frac{1}{|F|}\sum_{f\in F}h(_{t,f}).$ (8)
@@
 $\alpha$ and $k$ are cost parameters, $\bar{a}_{t}$ is the magnitude of the intervention.
-We employ the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) (Deb et al., 2002) to rank the generated candidates based on these three objectives.
+We employ the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) Deb et al.
+(2002) to rank the generated candidates based on these three objectives.
 Instead of selecting a single weighted average, we identify the Pareto Frontier where no objective can be improved without degrading another.
@@
 We evaluate our framework on three distinct critical care tasks, covering different datasets, demographics, and action space types.
-Sepsis Treatment We utilize the Medical Information Mart for Intensive Care (MIMIC)-IV v3.1(Johnson et al., 2020) database to learn optimal Intravenous (IV) fluid and Vasopressor administration strategies for sepsis patients.
-We select adult patients (age $\geq$ 18) satisfying the Sepsis-3 criteria (Singer et al., 2016), and exclude patients with missing mortality, demographics, and interventions, extreme missingness (>>60% missing data) or short hospital stays (<<24 hours).
+Sepsis Treatment We utilize the Medical Information Mart for Intensive Care (MIMIC)-IV v3.1Johnson et al.
+(2020) database to learn optimal Intravenous (IV) fluid and Vasopressor administration strategies for sepsis patients.
+We select adult patients (age $\geq$ 18) satisfying the Sepsis-3 criteria Singer et al.
+(2016), and exclude patients with missing mortality, demographics, and interventions, extreme missingness (>>60% missing data) or short hospital stays (<<24 hours).
 We define a septic shock starting from 24 hours before diagnosis and ending at 48 hours post-diagnosis.
@@
 We discretize IV fluid volume and vasopressor dosage into 5 bins each, resulting in a combined action space of $5\times 5=25$ discrete actions.
-Mechanical Ventilation We use the multi-center eICU Collaborative Research Database (eICU-CRD)(Pollard et al., 2018) to optimize ventilator settings for patients with respiratory failure.
+Mechanical Ventilation We use the multi-center eICU Collaborative Research Database (eICU-CRD)Pollard et al.
+(2018) to optimize ventilator settings for patients with respiratory failure.
 The cohort includes adult patients who underwent invasive mechanical ventilation (MV) for at least 24 hours.
@@
 We define a discrete action set based on clinically relevant adjustments for each parameter, resulting in a composite discrete action space of 18 actions (PEEP: 2 levels, FiO2: 3 levels, Tidal Volume: 3 levels).
-Renal Replacement Therapy We apply our method to the AmsterdamUMCdb (Thoral et al., 2021) to optimize continuous renal replacement therapy (RRT) dosing.
+Renal Replacement Therapy We apply our method to the AmsterdamUMCdb Thoral et al.
+(2021) to optimize continuous renal replacement therapy (RRT) dosing.
 We select adult patients (age $\geq$ 18) who received CRRT during the ICU stay where the stay is no longer than 60 days.
@@
 All continuous state features are normalized to zero mean and unit variance to ensure training stability.
-For policy training, we use common safe offline RL methods, Batch-Constrained Q-Learning (BCQ)(Fujimoto et al., 2019) for discrete action space tasks (Sepsis and Ventilation), and Implicit Q-Learning (IQL)(Kostrikov et al., 2021) for continuous action space task (RRT).
+For policy training, we use common safe offline RL methods, Batch-Constrained Q-Learning (BCQ)Fujimoto et al.
+(2019) for discrete action space tasks (Sepsis and Ventilation), and Implicit Q-Learning (IQL)Kostrikov et al.
+(2021) for continuous action space task (RRT).
 4.2 Baselines
@@
 4.2.1 Heuristic Reward Design
-We implement three standard reward structures commonly used in clinical RL literature. These serve as the ”human-expert” baselines, representing varying degrees of clinical foresight:
+We implement three standard reward structures commonly used in clinical RL literature. These serve as the "human-expert" baselines, representing varying degrees of clinical foresight:
 Outcome Reward Model (ORM): A purely outcome-oriented function that assigns a non-zero reward only at the terminal step, +100+100 for survival (discharge) and 100-100 for mortality.
-The value of $\pm 100$ is used in both Sepsis(Komorowski et al., 2018) and Ventilation(Peine et al., 2021) tasks in previous literature.
+The value of $\pm 100$ is used in both SepsisKomorowski et al.
+(2018) and VentilationPeine et al.
+(2021) tasks in previous literature.
 Since there is no past work that utilized outcome reward for RRT task, we adopt the same setting of $\pm 100$ for RRT.
@@
 It incentivizes short-term stabilization without long-term survival awareness.
-We use the reward function fully based on SOFA score(Zhang et al., 2024c) for Sepsis.
-For Ventilation, though we note that there is one process reward based on ventilator free days(Yousuf et al., 2025), the reward calculated on our cohort is not dense enough for efficient policy training, we therefore adopt the intermediate reward from the OPRM(Liu et al., 2024).
+We use the reward function fully based on SOFA scoreZhang et al.
+(2024c) for Sepsis.
+For Ventilation, though we note that there is one process reward based on ventilator free daysYousuf et al.
+(2025), the reward calculated on our cohort is not dense enough for efficient policy training, we therefore adopt the intermediate reward from the OPRMLiu et al.
+(2024).
 This further indicates that a fixed reward function may not be transferrable between different cohorts even with the same dataset.
-For RRT, we use the reward function based on Estimated Glomerular Filtration Rate (eGFR)(Zhang et al., 2024a).
+For RRT, we use the reward function based on Estimated Glomerular Filtration Rate (eGFR)Zhang et al.
+(2024a).
 Outcome Process Reward Model(OPRM): A weighted combination of the sparse and dense signals.
 This represents the state-of-the-art manual design, aiming to balance immediate physiological regularization with the ultimate goal of patient survival.
-We select from current SOTA methods (Perera et al., 2026)for Sepsis and (Liu et al., 2024)for Ventilation.
+We select from current SOTA methods Perera et al.
+(2026)for Sepsis and Liu et al.
+(2024)for Ventilation.
 Due to the lack of work on using OPRM for RRT, we choose to combine the PRM(eGFR) with ORM($\pm 100$) as OPRM.
@@
 4.5 RQ2: Can we use naive shaping of outcome reward model?
-To facilitate policy training previous literature actively utilizes clinically-guided dense rewards that include both terminal and intermediate rewards to form an OPRM(Raghu et al., 2017; Lu et al., 2024), which can be approximated as a naive reward shaping in a broader sense.
+To facilitate policy training previous literature actively utilizes clinically-guided dense rewards that include both terminal and intermediate rewards to form an OPRMRaghu et al.
+(2017); Lu et al.
+(2024), which can be approximated as a naive reward shaping in a broader sense.
 As shown in Figure2, the standalone process reward is not well correlated with mortality, which is dangerous as a secondary outcome to guide policy learning.
@@
 LLMs are rapidly developing and demonstrating excellent reasoning capabilities.
-Previous research has explored the possibility of directly using LLMs as planners for DTRs(Luo and Zhu, 2025) and concludes that they exhibit notable limitations.
+Previous research has explored the possibility of directly using LLMs as planners for DTRsLuo and Zhu (2025) and concludes that they exhibit notable limitations.
 Therefore, a natural question is whether they can instead directly serve as the reward signal instead and output a scalar quality score.
@@
 Appendix A Prompts
-Prompt 1: Prompt for Feature Selection
+Listing 1: Prompt for Feature Selection
 You are an expert Clinical Data Scientist and Intensivist specializing in Offline Reinforcement Learning.
@@
 - feature\_name: r= (p=, n=)
-Prompt 2: Prompt for Reward Generation
+Listing 2: Prompt for Reward Generation
 […]
@@
 80 """
-81 Strategic annealing time decay function.
-82 Exponential decay with half-life 48 hours.
-83 """
+81␣␣␣␣Strategicannealingtimedecayfunction.
+82␣␣␣␣Exponentialdecaywithhalf-life48hours.
+83␣␣␣␣"""
 84 half_life = 48.0
@@
 89 """
-90 Compute survival score for a single normalized feature value in [0,1].
-91 Types:
-92 - ’bell’: Gaussian bell curve centered at target with sigma.
-93 - ’decay_lower’: exponential decay if val > target.
-94 - ’directional_decay’: exponential decay from threshold or zero depending on direction.
-95 Returns score in [0,1].
-96 """
+90␣␣␣␣Computesurvivalscoreforasinglenormalizedfeaturevaluein[0,1].
+91␣␣␣␣Types:
+92␣␣␣␣-’bell’:Gaussianbellcurvecenteredattargetwithsigma.
+93␣␣␣␣-’decay_lower’:exponentialdecayifval>target.
+94␣␣␣␣-’directional_decay’:exponentialdecayfromthresholdorzerodependingondirection.
+95␣␣␣␣Returnsscorein[0,1].
+96␣␣␣␣"""
 97 if val is None:
@@
 148 """
-149 Exponential decay of confidence with delta_t (hours).
-150 Returns weight in (0,1].
-151 """
+149␣␣␣␣Exponentialdecayofconfidencewithdelta_t(hours).
+150␣␣␣␣Returnsweightin(0,1].
+151␣␣␣␣"""
 152 if delta_t is None or delta_t < 0:
@@
 157 """
-158 Compute competence penalty from actions.
-159 Penalize higher doses linearly scaled.
-160 Sum penalties for vaso and iv fluid, capped at 1.
-161 Returns penalty in [0,1].
-162 """
+158␣␣␣␣Computecompetencepenaltyfromactions.
+159␣␣␣␣Penalizehigherdoseslinearlyscaled.
+160␣␣␣␣Sumpenaltiesforvasoandivfluid,cappedat1.
+161␣␣␣␣Returnspenaltyin[0,1].
+162␣␣␣␣"""
 163 vaso_level = action.get(’vaso_5quantile’, 0)
@@
 174 """
-175 Compute potential function Phi(s,a,t) as balanced weighted sum of survival,
-176 confidence, and competence components with strategic time decay.
-177 state: dict of feature: (value_normalized, delta_t)
-178 action: dict with keys [’vaso_5quantile’, ’iv_fluid_5quantile’]
-179 t: absolute time step (int)
-180 """
+175␣␣␣␣ComputepotentialfunctionPhi(s,a,t)asbalancedweightedsumofsurvival,
+176␣␣␣␣confidence,andcompetencecomponentswithstrategictimedecay.
+177␣␣␣␣state:dictoffeature:(value_normalized,delta_t)
+178␣␣␣␣action:dictwithkeys[’vaso_5quantile’,’iv_fluid_5quantile’]
+179␣␣␣␣t:absolutetimestep(int)
+180␣␣␣␣"""
 181 survival_scores = []
@@
 224 """
-225 Reward is difference-based with discount factor:
-226 R(s,t,s’,t’) = gamma * Phi(s’,t’) - Phi(s,t)
-227 """
+225␣␣␣␣Rewardisdifference-basedwithdiscountfactor:
+226␣␣␣␣R(s,t,s’,t’)=gamma*Phi(s’,t’)-Phi(s,t)
+227␣␣␣␣"""
 228 return gamma * potential_function(s_next, t_next) - potential_function(s, t) - compute_competence_cost(a)
@@
 126 """
-127 Strategic annealing decay: potential decays toward zero as episode progresses.
-128 Use exponential decay with half-life of 48 hours (2 days).
-129 """
+127␣␣␣␣Strategicannealingdecay:potentialdecaystowardzeroasepisodeprogresses.
+128␣␣␣␣Useexponentialdecaywithhalf-lifeof48hours(2days).
+129␣␣␣␣"""
 130 half_life = 48.0
@@
 135 """
-136 Compute survival score for one feature normalized val in [0,1].
-137 params: dict with keys depending on type:
-138 - ’type’: ’bell’, ’decay_low’, ’decay_high’
-139 - For bell: ’target’, ’sigma’
-140 - For decay: ’tau’
-141 Returns score in [0,1].
-142 """
+136␣␣␣␣Computesurvivalscoreforonefeaturenormalizedvalin[0,1].
+137␣␣␣␣params:dictwithkeysdependingontype:
+138␣␣␣␣␣␣-’type’:’bell’,’decay_low’,’decay_high’
+139␣␣␣␣␣␣-Forbell:’target’,’sigma’
+140␣␣␣␣␣␣-Fordecay:’tau’
+141␣␣␣␣Returnsscorein[0,1].
+142␣␣␣␣"""
 143 if params[’type’] == ’bell’:
@@
 163 """
-164 Exponential decay of confidence with delta_t (hours).
-165 """
+164␣␣␣␣Exponentialdecayofconfidencewithdelta_t(hours).
+165␣␣␣␣"""
 166 return math.exp(-delta_t / tau)
@@
 169 """
-170 Compute penalty cost for actions.
-171 Higher dose (level) => higher penalty.
-172 Normalize each action level by max level, sum and scale.
-173 Return cost in [0,1].
-174 """
+170␣␣␣␣Computepenaltycostforactions.
+171␣␣␣␣Higherdose(level)=>higherpenalty.
+172␣␣␣␣Normalizeeachactionlevelbymaxlevel,sumandscale.
+173␣␣␣␣Returncostin[0,1].
+174␣␣␣␣"""
 175 total_cost = 0.0
@@
 188 """
-189 Compute potential function Phi(s,t) = survival * confidence
-190 with strategic time decay.
-191 state: dict of feature -> (val, delta_t)
-192 action: dict of action_name -> level
-193 t: absolute time step (int)
-194 """
+189␣␣␣␣ComputepotentialfunctionPhi(s,t)=survival*confidence
+190␣␣␣␣withstrategictimedecay.
+191␣␣␣␣state:dictoffeature->(val,delta_t)
+192␣␣␣␣action:dictofaction_name->level
+193␣␣␣␣t:absolutetimestep(int)
+194␣␣␣␣"""
 195 survival_scores = []
@@
 225 """
-226 Reward = gamma * Phi(s’,t’) - Phi(s,t)
-227 """
+226␣␣␣␣Reward=gamma*Phi(s’,t’)-Phi(s,t)
+227␣␣␣␣"""
 228 return gamma * potential_function(s_next, t_next) - potential_function(s, t) - compute_competence_cost(a)
@@
 106 """
-107 Strategic annealing decay: potential decays toward zero as episode progresses.
-108 Use exponential decay with half-life of 100 time steps.
-109 """
+107␣␣␣␣Strategicannealingdecay:potentialdecaystowardzeroasepisodeprogresses.
+108␣␣␣␣Useexponentialdecaywithhalf-lifeof100timesteps.
+109␣␣␣␣"""
 110 half_life = 100
@@
 115 """
-116 Compute survival score for one feature value normalized 0-1.
-117 For bell curve: Gaussian centered at target with sigma.
-118 For directional: exponential decay from threshold.
-119 Returns score in [0,1].
-120 """
+116␣␣␣␣Computesurvivalscoreforonefeaturevaluenormalized0-1.
+117␣␣␣␣Forbellcurve:Gaussiancenteredattargetwithsigma.
+118␣␣␣␣Fordirectional:exponentialdecayfromthreshold.
+119␣␣␣␣Returnsscorein[0,1].
+120␣␣␣␣"""
 121 if params[’type’] == ’bell’:
@@
 151 """
-152 Exponential decay of confidence with delta_t (hours).
-153 """
+152␣␣␣␣Exponentialdecayofconfidencewithdelta_t(hours).
+153␣␣␣␣"""
 154 return math.exp(-delta_t / tau)
@@
 157 """
-158 Penalize higher dose actions.
-159 action is dict with key ’action’ and integer dose level 0-3.
-160 Returns penalty in [0, 1].
-161 """
+158␣␣␣␣Penalizehigherdoseactions.
+159␣␣␣␣actionisdictwithkey’action’andintegerdoselevel0-3.
+160␣␣␣␣Returnspenaltyin[0,1].
+161␣␣␣␣"""
 162 dose = action.get(’action’, 0)
@@
 172 """
-173 Compute potential function Phi(s,t) as weighted sum of survival * confidence - competence cost,
-174 then apply strategic time decay.
-175 """
+173␣␣␣␣ComputepotentialfunctionPhi(s,t)asweightedsumofsurvival*confidence-competencecost,
+174␣␣␣␣thenapplystrategictimedecay.
+175␣␣␣␣"""
 176 survival_sum = 0.0
@@
 215 """
-216 Reward is difference-based with discount factor:
-217 R = gamma * Phi(s’,t’) - Phi(s,t)
-218 """
+216␣␣␣␣Rewardisdifference-basedwithdiscountfactor:
+217␣␣␣␣R=gamma*Phi(s’,t’)-Phi(s,t)
+218␣␣␣␣"""
 219 phi_next = potential_function(s_next, t_next)
@@
 D.1 Breaking Policy Invariance
-Proposition D.1 (Cost Regularization Breaks Invariance and Enforces Efficiency).
+Proposition 1 (Cost Regularization Breaks Invariance and Enforces Efficiency).
 Let the reward function be defined as $R(s_{t},a_{t},s_{t+1})=F(s_{t},s_{t+1})-\lambda\mathcal{C}(a_{t})$, where $F(s_{t},s_{t+1})=\gamma\Phi(s_{t+1})-\Phi(s_{t})$ is a potential-based shaping function and $\mathcal{C}(a_{t})>0$ is a strictly positive action cost.
@@
 Here, the first term depends only on the boundary states and the trajectory duration $T$.
-The factor $\gamma^{T}$ provides a strategic incentive to reduce $T$ (avoiding ”holding” behavior), but remains invariant to the path taken to reach $s_{T}$ at time $T$.
+The factor $\gamma^{T}$ provides a strategic incentive to reduce $T$ (avoiding "holding" behavior), but remains invariant to the path taken to reach $s_{T}$ at time $T$.
 However, the second term, $\lambda\sum\gamma^{t}\mathcal{C}(a_{t})$, strictly does not telescope.
@@
 D.2 Equivalence to Lagrangian Relaxation of CMDP
-Proposition D.2 (Equivalence to Lagrangian Relaxation of CMDP).
+Proposition 2 (Equivalence to Lagrangian Relaxation of CMDP).
 Maximizing the expected cumulative return of the regularized reward function $R(s,a,s^{\prime})=(\gamma\Phi(s^{\prime})-\Phi(s))-\lambda\mathcal{C}(a)$ is equivalent to maximizing the Lagrangian relaxation of a Constrained Markov Decision Process (CMDP), where the objective is to maximize physiological gain $J_{\Phi}$ subject to a constraint on the cumulative intervention cost $J_{\mathcal{C}}\leq\beta$.
@@
 Since the baseline approach involves a simple, single-pass inference to generate a policy, the computational cost is manageable even with expensive APIs.
-This ensures our method is compared against the strongest possible ”zero-shot” clinical agents, establishing a rigorous upper bound for performance without regarding training constraints.
+This ensures our method is compared against the strongest possible "zero-shot" clinical agents, establishing a rigorous upper bound for performance without regarding training constraints.
 In contrast, our medR framework requires generating dense reward signals across vast state-action spaces, making the use of commercial APIs prohibitively expensive and computationally intractable for large-scale training.
@@
 Traditional RL methodologies in healthcare typically define reward functions using clinical indicators, such as changes in SOFA scores, expert-derived heuristics, or terminal outcomes, including 90-day mortality.
-A foundational example is the AI Clinician by (Komorowski et al., 2018), which employed sparse, outcome-based rewards reflecting 90-day survival.
-A subsequent analysis(Huang et al., 2022) introduced safety constraints by limiting actions to those frequently taken by clinicians, further restricting exploration.
-(Raghu et al., 2017) addressed the limitations of sparse signals by incorporating intermediate clinical metrics such as SOFA, lactate, and increasing feedback frequency through their heuristic-based approach, which risks encoding clinical bias.
-(Raghu et al., 2018) and (Peng et al., 2018) advanced this direction by using model-based shaping, where predicted mortality log-odds serve as reward signals.
+A foundational example is the AI Clinician by [22], which employed sparse, outcome-based rewards reflecting 90-day survival.
+A subsequent analysis[17] introduced safety constraints by limiting actions to those frequently taken by clinicians, further restricting exploration.
+[41] addressed the limitations of sparse signals by incorporating intermediate clinical metrics such as SOFA, lactate, and increasing feedback frequency through their heuristic-based approach, which risks encoding clinical bias.
+[42] and [35] advanced this direction by using model-based shaping, where predicted mortality log-odds serve as reward signals.
 While this enhances continuity, it introduces dependence on opaque, learned predictors.
-(Futoma et al., 2017) employed high-fidelity simulators to define reward functions, though such strategies are resource-intensive and may lack clinical transparency.
+[14] employed high-fidelity simulators to define reward functions, though such strategies are resource-intensive and may lack clinical transparency.
 Despite these advances, most reward functions rely on proxy measures that lack semantic clarity and deep clinical grounding.
@@
 F.2 LLMs in Clinical Settings
-LLMs have transformed natural language tasks in clinical medicine (Thirunavukarasu et al., 2023; Haltaufderheide and Ranisch, 2024; Shool et al., 2025), aiding in medical question answering (Singhal et al., 2025), policy critique (Yang et al., 2025), and patient summaries(Asgari et al., 2025).
+LLMs have transformed natural language tasks in clinical medicine [52, 16, 46], aiding in medical question answering [50], policy critique [62], and patient summaries[1].
 However, their application in structured reward modelling for healthcare RL remains largely unexplored, presenting a promising opportunity.
 Recent studies have started exploring how LLMs can enhance RL in healthcare by improving sample efficiency and reward design.
-For example,(Bhambri et al., 2024) uses LLMs for abstract plans in potential-based reward engineering, boosting sample efficiency in toy environments, but their method relies on handcrafted abstractions and external verifiers.(Wu, 2024) proposes Q-engineering, which utilises LLM heuristics to shape Q-values, thereby improving efficiency but relying on task-specific heuristics and lacking interpretability.
-(Zhang et al., 2024b) shows how LLMs refine intrinsic rewards through code synthesis in rule-based environments, but their method is constrained by fixed domain rules and static strategies.
+For example,[4] uses LLMs for abstract plans in potential-based reward engineering, boosting sample efficiency in toy environments, but their method relies on handcrafted abstractions and external verifiers.[59] proposes Q-engineering, which utilises LLM heuristics to shape Q-values, thereby improving efficiency but relying on task-specific heuristics and lacking interpretability.
+[65] shows how LLMs refine intrinsic rewards through code synthesis in rule-based environments, but their method is constrained by fixed domain rules and static strategies.
 These methods advance LLM-guided reward design; however, they remain limited by the use of symbolic models, heuristics, and predefined knowledge, which reduces their applicability in complex clinical settings.
-These limitations are highlighted by (Gao et al., 2024), who explore reward models but face reward hacking issues.(Pternea et al., 2024) proposes an RL/LLM taxonomy but doesn’t address real-world clinical challenges, while (Sun, 2023) reviews RLHF but struggles with sparse feedback in healthcare.
+These limitations are highlighted by [15], who explore reward models but face reward hacking issues.[39] proposes an RL/LLM taxonomy but doesn’t address real-world clinical challenges, while [51] reviews RLHF but struggles with sparse feedback in healthcare.
 In contrast, our approach leverages LLMs for temporal credit assignment in clinical settings, transforming sparse outcomes into clinically interpretable rewards from raw patient data, without relying on external verifiers or predefined rules, enabling scalable, adaptive learning for dynamic healthcare environments.
 Recent studies have made significant strides in LLM-based reward engineering for RL.
-(Nazir and Banerjee, 2025) use zero-shot LLMs to replace human feedback, improving sample efficiency.(Fu et al., 2025) introduces PAR, addressing reward hacking in RLHF.
-(Deng et al., 2025) proposes the LMGT framework to balance exploration and exploitation, while (Wang et al., 2025a) develop SPA-RL for stepwise reward attribution.
-(Zhang et al., 2024b) proposes a simple framework using LLMs for intrinsic reward generation.
+[32] use zero-shot LLMs to replace human feedback, improving sample efficiency.[12] introduces PAR, addressing reward hacking in RLHF.
+[8] proposes the LMGT framework to balance exploration and exploitation, while [54] develop SPA-RL for stepwise reward attribution.
+[65] proposes a simple framework using LLMs for intrinsic reward generation.
 Despite these advancements, our approach is unique.
@@
 Temporal credit assignment remains a major challenge in RL, for complex, long-horizon tasks.
-Traditional methods, such as TD-lambda ((Sun, 2023)) and GAE ((Schulman et al., 2015)), focus on temporal differences but lack semantic reasoning, which limits their performance.
-Recent advances, such as Universal GAE by (Kwiatkowski et al., 2023), improve generalisation.
-Meanwhile, transformer-based TD learning by (Wang et al., 2025b), which integrates attention-based models, addresses some limitations but still falls short in complex, real-world tasks.
+Traditional methods, such as TD-lambda ([51]) and GAE ([45]), focus on temporal differences but lack semantic reasoning, which limits their performance.
+Recent advances, such as Universal GAE by [24], improve generalisation.
+Meanwhile, transformer-based TD learning by [55], which integrates attention-based models, addresses some limitations but still falls short in complex, real-world tasks.
+In addition, some works [58, 61, 7] attempt to analyze the timing of action provision from a causal perspective.
 LLMs have emerged as powerful solutions, bringing new levels of contextual understanding and reasoning to RL, effectively tackling long-standing challenges in temporal credit assignments.
-For example, (Pignatelli et al., 2024) introduces CALM, where LLMs act as critics to decompose tasks into subgoals and assign rewards in zero-shot settings; (Jiang et al., 2025) presents QLLM, replacing complex mixing networks in multi-agent RL with LLM-generated, interpretable credit functions; and (Qu et al., 2025) propose LaRe, using LLMs to derive multi-dimensional latent rewards that capture nuanced contributions in episodic tasks.
+For example, [37] introduces CALM, where LLMs act as critics to decompose tasks into subgoals and assign rewards in zero-shot settings; [19] presents QLLM, replacing complex mixing networks in multi-agent RL with LLM-generated, interpretable credit functions; and [40] propose LaRe, using LLMs to derive multi-dimensional latent rewards that capture nuanced contributions in episodic tasks.
 Collectively, these works establish LLMs as scalable, interpretable, and powerful tools for automating credit assignments in both single- and multi-agent environments.
 Our approach, medR, uniquely advances this line by leveraging real clinical data to integrate Large Language Models with potential-based reward shaping.
-Unlike CALM’s reliance on toy subgoals (Pignatelli et al., 2024), QLLM’s multi-agent focus (Pignatelli et al., 2024), or LaRe’s (Qu et al., 2025) latent outcome alignment, medR transforms sparse terminal outcomes into dense, mathematically grounded learning signals.
+Unlike CALM’s reliance on toy subgoals [37], QLLM’s multi-agent focus [37], or LaRe’s [40] latent outcome alignment, medR transforms sparse terminal outcomes into dense, mathematically grounded learning signals.
 By prompting LLMs to estimate the physiological stability of patient states ($\Phi(s)$), we construct a regularized reward structure that explicitly balances clinical improvement with intervention cost.